### v1

# Discretion in the Loop: Human Expertise in Algorithm-Assisted College Advising

Abstract
In higher education, many institutions use algorithmic alerts to flag at-risk students and deliver advising at scale. While much research has focused on evaluating algorithmic predictions, relatively little is known about how discretionary interventions by human experts shape outcomes in algorithm-assisted settings. We study this question using rich quantitative and qualitative data from a randomized controlled trial of an algorithm-assisted advising program at Georgia State University. Taking a mixed-methods approach, we examine whether and how advisors use context unavailable to an algorithm to guide interventions and influence student success. We develop a causal graphical framework for human expertise in the interventional setting, extending prior work on discretion in purely predictive settings. We then test a necessary condition for discretionary expertise using structured advisor logs and student outcomes data, identifying several interventions that meet the criterion for statistical significance. Accordingly, we estimate that 2 out of 3 interventions taken by advisors in the treatment arm were plausibly “expertly targeted” to students using non-algorithmic context. Systematic qualitative analysis of advisor notes corroborates these findings, showing that advisors incorporate diverse forms of contextual information—such as personal circumstances, financial issues, and student engagement—into their decisions. Finally, we explore the broader implications of human discretion for long-term outcomes and equity, using heterogeneous treatment effect estimation. Our results offer theoretical and practical insight into the real-world effectiveness of algorithm-supported college advising, and underscore the importance of accounting for human expertise in the design, evaluation, and implementation of algorithmic decision systems.
1 Introduction
As algorithm–assisted decision making becomes increasingly common in high-stakes public domains, there is growing recognition that evaluating these systems requires more than assessing model performance in isolation. In settings where algorithms inform—but do not replace—human decision-makers, understanding how human experts actually use algorithmic tools is essential to assessing their real-world impact and equity implications [29, 36, 59, 18]. For example, in college advising, an algorithm may identify students at risk of dropping out, but it is up to a human counselor to have a conversation with the student and suggest a course of action based on the student’s specific circumstances.
Evaluating human decision making when assisted by algorithms is essential for the design, evaluation, and implementation of automated decision systems. When algorithms are embedded in human decision processes, it becomes critical to ask: How much of the decision-making success is attributable to the algorithm versus the human expert? What discretionary work do humans perform that the algorithm cannot? What additional, unobserved information are important for improving outcomes? These questions help reveal whether key contextual or non-algorithmic information needs to be explicitly collected by humans, or whether algorithmic systems should be redesigned to incorporate this information.
Understanding the balance between algorithmic and human contributions becomes crucial for expanding and generalizing the educational intervention to new contexts (e.g., a new insitution). While algorithms can be deployed at scale without capacity constraints, the human component may introduce scaling challenges (e.g., social science interventions that perform well in initial studies but do not perform well at scale) due to changes in the implementation, resource constraints[13], or staff training [37, 46]. If human discretion plays a significant role in improving outcomes, then expanding the intervention requires either increasing the capacity of the existing experts or recruiting similarly qualified personnel.
We explore these questions in the context of college advising, a setting in which human decisions are both consequential and resource-constrained. State universities face growing pressure to improve student success and close equity gaps, and algorithm–assisted advising has been proposed as one strategy to help institutions scale proactive support [62, 24]. Yet the real-world impact of such systems depends crucially on how advisors use them.
There has long been interest in how predictive algorithms influence student outcomes, with early work focusing on identifying at-risk students through early warning systems at secondary and tertiary levels [43, 10, 66, 22, 51]. More recent work questioned the actionability and real-world impact of these systems [47, 48], with observational studies finding weak or null effects [21, 56]. Empirical studies of the impact of predictive systems are often limited by challenges in data collection and study implementation. Analyzing the causal effect of an algorithm–assisted system requires (1) longitudinal data on predictions, interventions, and downstream outcomes, which are not often documented at the same time, and (2) randomization of treatment assignment, which can be difficult to implement in educational institutions.
Additionally, quantitative data alone often fail to capture how practitioners interpret and act on algorithmic outputs. Quantification can abstract away the social context and judgment that shape real-world decisions [71]. Qualitative data—such as advisors’ meeting notes or child welfare caseworker casenotes—can provide insight into “street-level” discretionary work [69, 70] that is not reflected in structured data. Such data are rarely collected in tandem with quantitative measures, especially in education.
Our study leverages a unique opportunity to fill these gaps: we combine rich quantitative and qualitative data from a randomized controlled trial to investigate the role of human discretion in algorithm–assisted college advising. Specifically, we draw on data from the landmark MAAPS experiment at Georgia State University [2, 64, 65]—a multi-year randomized controlled trial of a data-driven educational intervention described as “intensive, proactive, technology-enhanced advisement and degree planning.” The goal of the MAAPS study was to identify at-risk students and administer timely interventions to support these students through both regular advising meetings and algorithmic early warning systems.
The MAAPS data analyzed in this paper includes comprehensive quantitative records of timestamped advisor interactions and student outcomes per semester. Additionally, advisors in the MAAPS program systematically logged interventions performed during their meetings with students (e.g., recommending courses or degree planning) and recorded qualitative comments on their conversation with the student and specific guidance. These qualitative comments, in conjunction with the quantitative outcomes, provide a unique window into human decision making alongside algorithmic tools.
Using a novel mixed-methods approach combining theoretical modeling, statistical testing, and qualitative analysis, we investigate how advisors leverage non-algorithmic information in student interventions in the MAAPS experiment, and explore the broader implications of discretionary decision making in algorithm-assisted advising. In meetings with students, advisors learn information about students’ interests, struggles, and personalities, creating an information asymmetry [60] between human decision makers and algorithmic tools that are entirely reliant on administrative data. Advisors may use this non-algorithmically available context to expertly target their interventions to students, improving performance over any possible algorithm. Our central research question is: What is the evidence for advisor discretion in algorithm–assisted advising, and what is its role in improving student outcomes?
Our contributions are as follows:
1. Theoretical framework for discretionary expertise in data-driven interventions. We develop a causal graphical framework to define and audit for human expertise in the interventional setting. Extending prior work on expertise in pure prediction tasks [5, 6], we define expertly targeted action in terms of three causal conditions (effectiveness, targeting, and heterogeneoty) and derive a statistical condition that is necessary but insufficient for proving expertise in interventions. We highlight a fundamental impossibility result in the interventional setting: discretionary expertise cannot be identified from observational data alone.
2. Empirical evidence of expert discretion in data-driven advising. Using quantitative data from the MAAPS randomized trial, we test for discretionary expertise by examining whether specific advisor interventions predict student outcomes (persistence, GPA) conditional on algorithmically available features. We identify five such ‘expert’ interventions ($p<0.05$) through a conditional independence test of the statistical condition identified in our theoretical framework. Subsequent qualitative analysis of advisor comments show that advisors indeed incorporate non-algorithmic context, such as the student’s personal circumstances, or financial troubles, into the ‘expert’ interventions.
3. Impact of advising on student outcomes and equity. We explore differences in advising styles across advisors, finding qualitative evidence of variation (i.e., academic vs. holistic focus) but no statistically significant advisor-specific effects on outcomes. A power analysis suggests our study is underpowered to detect many forms of heterogeneity, underscoring the methodological challenges of equity evaluation in real-world deployments.
These findings illustrate the critical role of human discretion in algorithm-assisted interventions and highlight the importance of human factors to the success of data-driven college advising.
2 Related Work
In recent years, there has been a growing interest in the downstream effects of algorithm-assisted human decisions, particularly within criminal justice, economics, healthcare, and higher education [49]. We review key contributions in these areas, contextualizing how they inform our study of algorithm-assisted advising.
Causal Inference in Algorithmic Decision Making.
Causal inference techniques are critical for evaluating algorithm-assisted decisions and their downstream effects. Notably, Imai et al. [36] conducted the first randomized field experiment to evaluate how an algorithmic decision aid in pre-trial risk assessment affects human judgment and decision quality in the U.S. criminal justice system. Follow-up work by Ben-Michael et al. [12] focused on whether algorithm-assisted human decisions improved decision-making “accuracy” over counterfactual baselines. Our work shifts the focus from decision quality per se to the mechanisms of human discretion: we investigate whether, how, and when human decision-makers contribute additional, outcome-relevant expertise. This is especially relevant in the educational setting where advisors choose among multiple possible interventions that may improve student outcomes.
In the educational domain, while the importance of understanding downstream effects of algorithmic student risk assessment has long been recognized [10], there has been little randomized evaluation of algorithmic decision making, aside from the current work, to the best of our knowledge. Perdomo et al. [55] use a regression discontinuity design to estimate the causal effect of early warning system risk scores on student success in Wisconsin public schools, finding small effects but have limited insight into underlying mechanisms due to missing data on subsequent interventions. Liu et al. [47] introduce a causal framework showing that accurate outcome prediction is insufficient for selecting effective interventions to improve individual outcomes in education.
Green and Chen [29] examine how algorithmic risk scores influence human decision making in a lab study, finding that human evaluators exhibit selective adherence to algorithmic recommendations, often amplifying existing biases. Raji and Liu [59] discuss methodological challenges of evaluating algorithmic-assisted decisions, emphasizing the need for experimental designs that disentangle the potential cognitive biases of a human judge.
Bias, Expertise, and Discretion in Education and Related Domains.
The role of evaluator expertise and bias in algorithm-assisted decisions has been a central concern in multiple domains. Holstein and Doroudi [33] and Baker and Hawn [8] explore how biases in algorithmic predictions can propagate through human decision making, particularly in educational contexts. Feathers [23] also examines how race and other demographic factors influence algorithmic predictions in higher education, raising equity and fairness concerns. These findings align with broader debates on the unintended consequences of algorithmic interventions in advising and student support.
Li [45] further investigates how expertise in peer review influences evaluators’ reliance on algorithmic recommendations, showing that domain experts may either mitigate or exacerbate algorithmic biases depending on their pre-existing heuristics and decision-making frameworks. Likewise, Hoffman et al. [32] analyze how hiring managers integrate algorithmic predictions into their decisions, demonstrating that discretion can either counteract or reinforce algorithmic biases. Gillis et al. [28] show that whether the algorithm amplifies or reduces the bias of the human decision-maker, with respect to a protected class, depends on whether the algorithm uses the protected class as a feature.
Information Asymmetries and Reliance on Algorithms.
Even in settings where algorithms outperform human decision-makers, individuals may possess private information that could refine algorithmic recommendations. Our quantitative approach to studying these information asymmetries is most similar to Alur et al. [5]’s. They devise a statistical test to audit for human use of “algorithmically-unavailable information” in machine-assisted prediction. We apply their test to the MAAPS dataset with slight modifications (described in §5.1). Overall, our work differs significantly with four key contributions. First, we generalize their findings to settings where human experts use non-algorithmic information to take actions rather than make predictions. Second, we formalize the expert decision problem using structural causal modeling and prove identifiability results (§4). Third, we develop a novel mixed-methods analytical framework to complement the quantitative test (§3.2), coding advisor comments to identify the particular non-algorithmic information guiding advisor decision making. Fourth, we apply our methods to a distinct domain—educational interventions in the MAAPS study—whereas Alur et al. [5] study medical triage.
In addition to Alur et al. [5], a growing body of work studies how humans combine their own judgment with the algorithmic predictions and what scenarios lead to humans relying more on algorithms. Balakrishnan et al. [9] show that people often combine their judgments with algorithmic recommendations, increasing prediction errors by over-relying on algorithms when their own insights are more valuable and vice versa. Snyder et al. [73] find that humans rely on algorithms when the algorithms perform better than humans (on average) and when the humans are required to make decisions quickly.
Recent work by Hullman et al. [35] developed a decision-theoretic framework to formalize when human decision-makers can be expected to make utility-maximizing choices in algorithm-assisted contexts, and applied it to quantify the value of information provided by such systems in synthetic decision-making tasks [31]. In contrast, our model of advisor decision making does not assume a rational, utility-maximizing agent, and our goals are to empirically assess whether and how advisors non-algorithmic information in a real world deployment of algorithm-assisted advising.
We discuss further broadly related work in §8.
3 Overview of Data and Methodology
This section provides an overview of the data and methodology used in our study. In §3.1, we describe the MAAPS randomized control trial at Georgia State University, and the structured and unstructured data sources we analyze. §3.2 outlines our mixed-methods analytical framework, explaining how we integrate quantitative and qualitative approaches to study advisor discretion and expertise.
3.1 MAAPS Experiment and Data
3.1.1 Experiment
Our data comes from Georgia State University’s participation in a randomized control trial run by the National Institute for Student Success (NISS) at eleven institutions from 2016 to 2022. The trial tested for effects of an intensive advising protocol called the Monitoring Advising Analytics to Promote Success (MAAPS) intervention, which has three pillars: (1) a data-driven early warning system; (2) regular review of degree plans; and (3) administering targeted and timely interventions based on regular review and early alerts.
The MAAPS data driven early alert system (1) at GSU comprised of an academic risk predictor (“predicted supported level”) that was embedded within a third party software, which advisors could access but students could not. [^1] EAB’s Navigate student success management platform software. Historical risk prediction data and model details for the MAAPS cohort were unavailable for this study.
At GSU, first-year students are assigned advisors from a centralized office who provide them with academic guidance from enrollment to graduation. The MAAPS intervention was implemented among the entering class of 2016, assigning students in the treatment group to a select group of five specially trained advisors to implement the MAAPS protocol, while those in the control group were assigned to standard academic advisors. [^2] At any one time, there were three active MAAPS advisors. Due to advisor turnover, there were five total MAAPS advisors over the course of the experiment. In fall 2017, one of the MAAPS advisors (Advisor A) left GSU and was replaced by a successor who took over until summer 2018, after which this advisor left and was replaced by another.
At the four-year mark, students who received MAAPS advising demonstrated a statistically significant increase in GPA (by $0.16$ points on average) and credit accumulation (by 6 credits on average) [64]. At the six-year mark, students in the treatment group showed a statistically significant increase in graduation rate (by $7\%$) [65]. [^3] GSU was the only institution to show significant positive treatment effects of the MAAPS intervention; we discuss potential reasons for this in §8.1.
3.1.2 Data
We analyze data from the first four years of GSU’s participation in the study in order to understand advisor’s discretionary decision making in the treatment arm of the experiment.
Structured Student Baseline Features and Outcomes Data.
For all 1037 students (520 in treatment and 517 in control) who participated in the MAAPS study, the dataset includes their baseline characteristics at the time of enrollment to GSU (such as their demographics, high school GPA and their scores on college entry tests), as well as timeseries data on their educational outcomes updated each term (such as their GPA and credit accumulation), including their graduation outcomes. This data allows us to understand how discretionary decision making connects to students’ educational outcomes and to look for heterogeneity of treatment effects in their baseline features. A full list of dataset variables are in Appendix A.1.
Structured Data on Advisor Interventions.
We also rely on logs that advisors in the treatment arm kept on all their interactions with students, providing a detailed picture of their decision-making processes. The five advisors working with treatment students in the MAAPS experiment held 5,269 meetings with students. For each meeting, advisors recorded structured data on interventions they performed during the interaction (selected from a list of 20, in Appendix A.1.3, for instance advising students to correct their schedule, see their instructor, or seek career counseling). We use this data to audit quantitatively for actions taken by advisors that cannot be “explained” by algorithmic prediction.
Unstructured Advisor Meeting Comments.
The advisor logs also include 1,888 free-form comments, in which advisors optionally elaborate on meetings. We use these comments to perform a complementary qualitative analysis of advisors’ decision making.
3.1.3 Data Quality and Cleaning
During the data pre-processing and cleaning phase, we remove students with missing data for transfer status, high school GPA, and expected family contribution (EFC) for all quantitative analyses. [^4] Following exclusion of students missing data on transfer status, all students were non-transfers/first-time freshmen. For all tests run on outcomes derived from timeseries data, we exclude data from summer semesters due to lower student enrollment and different student behaviors over summer semesters as compared to the academic year. Finally, we perform test-specific exclusions, which are described in their respective methods sections.
3.2 Our Mixed-Methods Approach
This study follows a mixed-methods approach, integrating a theoretical framework of advisor decision making (§4) with both quantitative (§5) and qualitative data analysis (§6). As aforementioned, primary analysis of the MAAPS RCT [64, 65] focused on estimating average treatment effects at the population level and was purely quantitative. [^5] While detailed data on advisor meetings was collected as part of the experiment, it was not previously analyzed. This study directly addresses this gap while introducing new research questions.
Our mixed-methods analytical framework is designed to investigate the role of advisor discretion and expertise within the treatment group. Our motivation for integrating quantitative and qualitative methods is three-fold: (1) to develop a comprehensive understanding of advisor expertise in interventional decision making, where qualitative insights help interpret and contextualize quantitative findings; (2) to support an exploratory analysis of heterogeneous treatment effects, especially in cases where quantitative signals may be subtle or underpowered; and (3) to produce novel qualitative insights that can directly inform professional development for advisors, reveal gaps, and support policy recommendations.
Our mixed-methods analytical framework offers a generalizable approach for studying human discretion and expertise in other settings involving algorithmically assisted decision making. The combination of causal inference, statistical testing, and qualitative analysis provides a holistic set of tools that extend beyond college advising to domains such as healthcare, social services, and hiring, where understanding the interaction between human judgment and algorithmic support is equally critical.
We reflect below on our roles as researchers and the context in which we conducted this study. Further details on our methods are in the relevant sections of the paper.
Researcher positionality.
We begin by reflecting on our positionality as researchers in this study. We recognize that our identities, disciplinary training, and institutional affiliations shape how we approached the research questions, interpreted the data, and interacted with the broader context of advising and algorithmic decision making. The research team includes graduate students (Authors 1 and 2) and tenure-track faculty (Authors 5 and 6) at R1 institutions, with training in computer science, operations research, and the societal impacts of algorithmic decision making, including secondary expertise in educational applications.
The research process was shaped by an ongoing, collaborative partnership between the research team and staff at the National Institute for Student Success (NISS). From November 2023 to November 2024, Authors 5 and 6 held biweekly meetings with three partnering staff members at NISS, including Authors 3 and 4, to understand the MAAPS intervention and its implementation context, as well as to determine the structure and limitations of the advising data. These meetings informed both our quantitative and qualitative analyses by grounding them in the realities of the program.
Our access to the data was limited to de-identified records shared under a data use agreement with NISS, and we did not interact directly with students or advisors. The Princeton University and Columbia University Institutional Review Boards determined the research was exempt from further review, as it does not qualify as human subjects research. We recognize that our interpretation of advising practices is shaped by our outsider perspective and by the structure and limitations of the available data. In the qualitative analysis, we sought to preserve the nuance of advisors’ written reflections while acknowledging the constraints of interpreting them at a distance.
Finally, as this study performs a secondary analysis of an experiment designed primarily to estimate average treatment effects, the original experimental design and data availability have shaped both the scope of our research questions and methodological choices.
4 Theoretical Framework: Expertly Targeted Interventions
A student meets with their advisor, explaining that a medical condition prevents them from walking long distances. In response, the advisor works with the student to design a schedule with nearby classrooms that also meets her academic interests. The student is able to achieve a high semester GPA.
This example, based on a real case, illustrates how advisors can draw on rich, algorithmically “invisible” information to support students in achieving their goals—in this case, a health condition that impacts the student’s schedule. These insights create an information asymmetry between human decision makers and algorithmic tools, raising the question: what role does human discretion play in delivering interventions and improving outcomes?
In this section, we propose a causal framework to understand the role of human expertise in improving outcomes when algorithmic decision aides are deployed in interventional settings, such as the MAAPS collegiate advising case.
We begin in §4.1 by introducing our model of algorithm-assisted intervention, which encapsulates the model of Alur et al. [5] for prediction as a special case. §4.2 defines the concept of human expertise in targeted actions and specifies assumptions about the causal mechanism. Finally, §4.3 provides theoretical results on detecting the presence expertly targeted actions from data. We show that, without further assumptions, it is impossible to detect expertly targeted actions from observational quantitative data alone. However, we define statistical conditions that can provide important evidence suggestive of expertly targeted action.
4.1 Model of Expertise in Interventions
We begin by providing a causal model of algorithm-assisted decision-making processes, and a causal definition of expertise in this setting. We use the machinery of structural causal models (SCMs), which define the way that variables causally influence one another through a set of functional relationships [53, 54, 57, 11].
We consider a model where a human decision-maker, at each time step $t$, needs to make a decision—corresponding in the MAAPS study to the action taken by an advisor in student meetings.
Let $\mathbf{X}_{t}$ be the algorithmically-observed features measured at time $t$. In the MAAPS experiment, $\mathbf{X}_{t}$ refers to the observable administrative data on the student. Let $Y_{t+1}$ denote the outcome variable of interest at the next time step, for instance, the student’s term GPA. We define $\mathbf{U}_{t}$ as the context at time $t$ that is relevant to the outcome $Y_{t+1}$ but unobserved by the algorithm; for instance the information learned through advising conversations. Here, outcome relevance means that $\mathbf{U}_{t}$ is an input of the process generating the student’s outcomes $Y_{t+1}$, to be formalized shortly. [^6] If there is no such relevant context, then $\mathbf{U}_{t}$ may be excluded from $\mathbf{V}_{\mathcal{A}}$.
The advisor observes $\mathbf{X}_{t}$ and potentially $\mathbf{U}_{t}$ and then takes action $A_{t}$ (e.g., advising the student to talk to their instructor).
[…]
For each variable $V\in\mathbf{V}_{\mathcal{A}}$, we define a function $f_{V}\in\mathcal{F}_{\mathcal{A}}$ that models $V$ as the causal output of a subset of other variables in $\mathbf{V}_{\mathcal{A}}$, together with an exogenous noise variable $\epsilon_{V}\in\mathbf{\epsilon}_{\mathcal{A}}$ specific to $V$.
We assume that these functions satisfy the input–output constraints specified by the directed acyclic graph (DAG) in Figure 1, where a directed edge from variable $V_{1}$ to $V_{2}$ indicates that $f_{V_{2}}$ takes $V_{1}$ as an input. For example, the structural equation $Y_{t+1}=f_{Y}(\mathbf{X}_{t},\mathbf{U}_{t},A_{t})$ corresponds to a causal graph in which $\mathbf{X}_{t}$, $\mathbf{U}_{t}$, and $A_{t}$ are direct causes (parents) of $Y_{t+1}$. [^8] Note that when no endogenous context $\mathbf{U}_{t}$ is in the model, i.e. $\mathbf{U}_{t}\notin\mathbf{V}_{\mathcal{A}}$, then $\mathbf{U}_{t}$ will not be a causal parent of $Y_{t+1}$.
Our model generalizes the functional model of human expertise in prediction, presented in Alur et al. [5], in two ways. First, we fully specify the causal structure among variables using DAGs. Second, we move beyond prediction to intervention, accounting for the possibility that the human decision-maker’s choices can influence future outcomes. [^9] Specifically, we replace the prediction variable $\hat{Y}$ [see eq. 5, 5]—which was assumed to have no effect on the predictive target $Y$—by an action variable $A_{t}$. The action $A_{t}$ may have a causal effect on the future outcome $Y_{t+1}$. Also, the action $A_{t}$ may be informed by both an algorithmic signal (represented by $\mathbf{X}_{t}\to A_{t}$) and additional, unobserved contextual knowledge (represented by $\mathbf{U}_{t}\to A_{t}$). [^10] Our model can capture three types of algorithm-assisted decision making of increasing generality. In the special case $A_{t}$ is performing an idle prediction $\hat{Y}$, we can simply stipulate $A_{t}$ does not influence the outcome $Y_{t+1}$ and our model reduces to Alur et al. [5]’s model. But unlike Alur et al. 5, our model can also represent predictive decision making, where predictions are made of an outcome in order to influence that outcome. With $R(X)$ denoting the range of a random variable $X$, we may think of $A$ as determined by a (potentially noisy) decision policy $\pi:R(\hat{Y})\to R(A)$ on the human’s internal predictions $\hat{Y}$. Finally, our model can represent more general algorithm-assisted interventional contexts, in which actions to affect an outcome do not necessarily depend on predictions of that outcome. Collegiate advising is one such example: An advisor may decide to take actions to support students for many reasons that are not all well-modeled as predictions of the outcome they aim to influence.
[…]
Figure 1: Causal graphical model of a single-stage, algorithm-assisted interventional decision setting. Dotted edges represent potential functional dependencies.
Finally, the model specifies a joint distribution $\mathbb{P}$ over the exogenous noise variables $\mathbf{\epsilon_{\mathcal{A}}}$, assumed to be independent. Together with the structural equations, this distribution induces unique observational and causal distributions on all model variables. The observational distribution is Markovian with respect to the causal DAG. [^11] That is, every variable is independent of its non-descendants given its parents. As is standard for causal identification from observational data, we assume the model is causally minimal, that is, the distribution is not Markovian with respect to any proper subgraph [57, pp. 107-109].
4.2 Definition of Expertly Targeted Action
Say a MAAPS advisor chooses to recommend a particular course to a student in a meeting. Given a full causal model of the decision-making scenario, when does this course recommendation demonstrate expertise? Here are three intuitive requirements:
  1. 1.
    The recommendation helps the student’s outcomes, e.g. their GPA. For instance, it should not be that the student ends up performing worse than they would have without taking the course.
  2. 2.
    The recommendation was targeted based on outcome-relevant information unavailable to the decision support algorithm —such as insights from a conversation about the student’s interests. For example, it should not be that the advisor recommends this course to every student, or that the recommendation is determined by the algorithm’s output.
  3. 3.
    The recommendation is effective because of the targeting. For instance, it should not be that all students who take this course get an A+ (and so the targeting was unimportant to the effectiveness of the intervention), but that the course recommendation uniquely benefits the targeted student.
These requirements can be stated as three causal statements—one about the action’s effect, one about the action’s targeting, and one about their interaction. When the three are true together, we call that expertly targeted action. To simplify the statement of the definition, we consider a setting in which both the action $A_{t}$ and external information $U_{t}$ are binary variables. [^12] We discuss in Appendix A.2.3 how to generalize the definition to accommodate non-binary contrasts.
Definition 4.1 (Expertise).
For a given observation with $\mathbf{X}_{t}=\mathbf{x}$, $U_{t}=u$, and $A_{t}=1$, a human decisionmaker’s action is expertly targeted when the following three conditions are met:
  1. 1.
    Effective action. Taking the action improves the outcome.
    $\mathbb{P}(Y_{t+1}=1\mid do(A_{t}=1),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)>\mathbb{P}(Y_{t+1}=1\mid do(A_{t}=0),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)$
  2. 2.
    Targeted action. The human decisionmaker chose to take the action based on non-algorithmic information.
    $\mathbb{P}(A_{t}=1|do(U_{t}=u),\mathbf{X}_{t}=\mathbf{x})>\mathbb{P}(A_{t}=1|do(U_{t}=1-u),\mathbf{X}_{t}=\mathbf{x})$
  3. 3.
    Heterogeneous action. The relative effectiveness of the action is due to heterogeneity in $U$.
    $\mathbb{P}(Y_{t+1}=1|do(U_{t}=u,A_{t}=1),\mathbf{X}_{t}=\mathbf{x})$ $-\mathbb{P}(Y_{t+1}=1|do(U_{t}=u,A_{t}=0),\mathbf{X}_{t}=\mathbf{x})>$
    $\mathbb{P}(Y_{t+1}=1|do(U_{t}=1-u,A_{t}=1),\mathbf{X}_{t}=\mathbf{x})$ $-\mathbb{P}(Y_{t+1}=1|do(U_{t}=1-u,A_{t}=0),\mathbf{X}_{t}=\mathbf{x})$
To capture instances of expertise according to this definition, it is important to describe actions at the right level of abstraction. For instance, say the action, $A_{t}$, is defined as “recommending a schedule.” If advisors recommends all students a schedule, then the targeted action component of the expertise definition will fail—even if the advisors are incorporating non-algorithmic information into the particular schedules they recommend to students.
4.3 Impossibility of Identifying Expertise from (Quantitative) Observational Data
The above definition considers how we would identify instances of expertise with full access to the causal mechanisms. However, in most cases we will only have an observational dataset $\mathcal{D}$ of $n$ $(\mathbf{x_{i}},y_{i},\hat{y}_{i})$, observations drawn from the joint distribution over variables $\mathbf{X}_{t}$, $A_{t}$ and $Y_{t+1}$. [^13] This observational distribution excludes $U_{t}$ because it is uncommon to collect data on $U_{t}$, the private, algorithmically-invisible information that humans incorporate in their decision making (and moreover, it is unclear how to systematically collect it). If we do not make any causal assumptions beyond those in §4.1 and Figure 1, can we identify whether an advisor demonstrates expertise in targeting from this observational information?
We prove an impossibility result that we cannot detect human expertise from observational data alone, because causal models that differ on the relevant causal facts can be statistically indistinguishable. However, we show that a related statistical condition—conditional dependence of actions $A_{t}$ and future outcomes $Y_{t+1}$ given algorithmically available data $\mathbf{X}_{t}$—is a necessary condition for expertly targeted action, and a sufficient condition for weakened components of it. We then test whether this condition is met in the MAAPS dataset in §5.
4.3.1 Impossibility Result on Identifying Expertise from Observational Data
The following proposition formalizes the nonidentifiability of expertly targeted action from observational data alone. The proof presents a counterexample of a pair of observationally equivalent SCMs that differ in whether Definition 4.1 is met.
Proposition 4.1 (Non-identifiability of expertly targeted action).
Expertly targeted action is impossible to detect from facts about the observational distribution over $\mathbf{X}_{t}$, $A_{t}$ and $Y_{t+1}$ alone.
Proof Sketch.
To determine whether expertise is present, we must establish both that $A_{t}$ affects $Y_{t+1}$ and that they are related by a common cause $U_{t}$ (Figure 4). In general, it is impossible to distinguish whether two variables are directly related, related by a common cause, or both using just observational data about those two variables [57]. We carefully construct two SCMs $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ (see complete proof in Appendix A.2.1). These SCMs induce identical observational distributions $\mathbb{P}(A_{t},Y_{t+1})$. $\mathcal{M}_{1}$ meets the definition of an expertly targeted action, while $\mathcal{M}_{2}$ does not. We achieve observational equivalence by tuning the exogenous noise variables in both models to produce the same distribution over $A_{t}$ and $Y_{t+1}$, even though their causal structures differ. ∎
Since it is impossible to identify expertly targeted action from observational data alone, qualitative data on how human decision-makers choose actions is crucial to understanding the deployment dynamics of algorithmic decision support systems. In §6.1, we describe how to use advisors’ qualitative logs of student interactions to supplement our understanding of whether and how they use non-algorithmic information to target their actions.
4.3.2 Necessary Statistical Condition for Expertise
We now consider a conditional dependence condition $A_{t}\not\perp Y_{t+1}|\textbf{X}_{t}$ that is a necessary condition for expertise to be present (under a further assumption). This condition is analogous to the conditional dependence condition used by Alur et al. [5]. [^14] Alur et al. [5] identify this conditional dependence condition as sufficient for expertise in the predictive case, since they define expertise as simply the presence of a causal influence of non-algorithmic information on predictions. However, this definition of expertise does not make sense in the interventional case; we should expect advisors’ actions to influence the outcomes positively to count as expertly taken. To establish the necessity of the conditional independence condition for expertly targeted action, we make the assumption that the SCM $\mathcal{A}$ is causally faithful.
Definition 4.2 (Causal faithfulness).
An SCM $\mathcal{M}$ is causally faithful when for all disjoint subsets of variables $\mathbf{X},\mathbf{Y},\mathbf{Z}\subseteq\mathbf{V}_{\mathcal{M}}$, $\mathbf{X}\perp\mathbf{Y}|\mathbf{Z}\implies\mathbf{X}\perp_{\mathcal{G}_{\mathcal{M}}}\mathbf{Y}\ |\ \mathbf{Z}$, where $\perp_{\mathcal{G}_{\mathcal{M}}}$ is the relation of d-separation relative to the model’s induced causal graph $\mathcal{G}_{\mathcal{M}}$.
The causal faithfulness condition is a standard non-degeneracy assumption, commonly made in causal inference. It is violated when causal effects perfectly cancel – in our case, if $A_{t}\leftarrow\mathbf{U}_{t}\to Y_{t+1}$ perfectly offsets $A_{t}\to Y_{t+1}$ – which is quite rare. [^15] For instance for linear models, it happens with zero probability when coefficients are drawn randomly [57, 74].
Proposition 4.2 (Necessary condition for expertise).
Assume causal faithfulness. If there is expertly targeted action, then $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$.
Proof.
For any of the three components of the expertly targeted action definition (Definition 4.1) to be met, it must be the case that $A_{t}\not\perp_{\mathcal{G}_{\mathcal{M}}}Y_{t+1}|\ \mathbf{X}_{t}$ (as Effective action requires $A_{t}\rightarrow Y_{t+1}$, Targeted action requires $U_{t}\rightarrow A_{t}$, and Heterogeneous action requires both). So then by causal faithfulness, $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$. ∎
Proposition 4.2 shows that, under causal faithfulness, a necessary condition for expertly targeted action is conditional dependence of the advisor’s action $A_{t}$ and the student’s outcome $Y_{t+1}|\mathbf{X}_{t}$. This proposition suggests a valuable method—by detecting conditional dependence, we detect a superset of actions exhibiting human expertise (putting power issues aside). We adapt the method from Alur et al. [5] to conduct this test in §5 and identify interventions meeting this necessary condition for expertise.
4.3.3 Sufficient Statistical Condition for Weakened Components of Expertise
Given the non-identifiability of expertly targeted actions (Proposition 4.1), the conditional dependence of actions $A_{t}$ and future outcomes $Y_{t+1}$ given $\mathbf{X}_{t}$ cannot be both a necessary and sufficient condition for expertly targeted actions. However, while insufficient to guarantee expertise, the statistical criterion can guarantee non-trivial components of it are met. To state the result, we define two concepts that weaken key components of the expertly targeted action definition, beginning with effective action:
Definition 4.3.
An action is impactful if for some $\mathbf{X_{t}=x}$, $U_{t}=u$,
$\mathbb{P}(Y_{t+1}=1|do(A_{t}=1),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)\neq\mathbb{P}(Y_{t+1}=1|do(A_{t}=0),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)$
Impactful action guarantees that actions change the outcome for some individual, equivalent to the presence of edge $A_{t}\to Y_{t+1}$ (Figure 4.1). Effective actions, which improve the outcome, are a subset of impactful actions.
We can define a similar weakened criterion for targeting, guaranteeing that non-algorithmic context influences the advisor at some point in the decision-making process (Figure 4.2):
Definition 4.4.
An action is non-algorithmic if for some $\mathbf{X}_{t}=x$,
$\mathbb{P}(A_{t}=1|do(U_{t}=1),\mathbf{X}_{t}=\mathbf{x})\neq\mathbb{P}(A_{t}=1|do(U_{t}=0),\mathbf{X}_{t}=\mathbf{x})$
Similar to impactful actions vs. effective action, non-algorithmic actions are a superset of targeted actions.
The following proposition demonstrates that meeting the statistical conditional dependence condition provides suggestive evidence for targeting, by showing that at least one of these necessary conditions of targeting are met. We defer its proof to Appendix A.2.2.
Proposition 4.3.
If $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$, then at least one of impactful action (Definition 4.3) or non-algorithmic action (Definition 4.1.2) are met.
Thus, if the statistical condition is met by an observational dataset, we can be certain that two necessary conditions for expertise are met: Within the entire population, there are individuals for whom the advisor’s actions are having an impact, are influenced by non-algorithmic information, or both. On the one hand, this provides another justification for using the condition to identify possible instances of expertly targeted action, without requiring a faithfulness assumption (Definition 4.2). On the other hand, the characterization in Proposition 4.3 brings to light several key limitations in interpreting the conditional independence condition; notably it does not establish whether the effect of the action on the outcome is heterogeneous with respect to $U_{t}$. [^16] As aforementioned, the condition does not guarantee expert targeting is present. Further, there is no guarantee that impactful action and non-algorithmic targeting are both met, or that they are met for the same individuals identified by $\mathbf{X}_{t}$ and $U_{t}$. The condition also does not confirm whether the action improves the chance of a good outcome, rather than merely changing those chances.
4.3.4 Implications for Expertise Auditing
The theoretical framework suggests that testing for the conditional dependence of the actions $A_{t}$ and the outcomes $Y_{t+1}|\mathbf{X}_{t}$ allows us to identify interventions meeting a necessary condition of expertly targeted action, and a sufficient condition for weakened components of it. In the MAAPS dataset, this corresponds to testing for the conditional dependence between (1) advisors’ actions, as indicated by advisors in their meeting logs, and (2) students’ future outcomes, such as GPA or persistence, given (3) their algorithmically-available features. We show how to implement this test with our data in §5.
The theoretical framework also clarifies what is impossible to ascertain from quantitative observational data. In particular, since the condition is only sufficient for weakened components of expertly targeted action, the framework illuminates three important concerns to attend to in event of conditional dependence to determine whether expertly targeted action is genuinely present. These concerns guide our interpretation of our quantitative results (§5.2.1) as well as our qualitative analysis of advisor comments (§6).
(A) Non-Algorithmic Targeting.
Since the test only identifies a necessary condition for expertly targeted action, it is important to identify whether there is corroborating evidence that human decisionmakers draw on non-algorithmic information to guide their interventions, for instance using qualitative data about the decision-making process. This can substantiate the claim that $\mathbf{U}_{t}$, the non-algorithmic context, reflects non-algorithmic sources of information that advisors are incorporating into their interventions to improve students’ outcomes.
(B) Supporting At-Risk Students.
In many cases, it may be that advisors are targeting their supportive actions to students based on algorithmically-unobserved risk factors: pieces of information about the student that suggest they are at risk for poor outcomes but are not captured by features of risk prediction (for instance, an unpredictable family emergency). However, even if actions taken by advisors counterfactually help these students, it may appear as if they are correlated with negative outcomes because the actions are targeted to an especially at-risk population. Thus, identifying corroborating evidence for targeting based on risk factors can help disentangle two very different explanations of a test result suggesting such negative correlation: that actions are hurting students, or supporting at-risk students. [^17] It might be tempting to try to disentangle Impactful action (Definition 4.3) from Effective action (Definition 4.1.2) observationally by looking just for positive correlations between taking an action and an outcome. However, this strategy is untenable because it is always possible that especially at-risk students, even if helped through expertly targeted action, may perform worse than average, leading to negative correlation.
(C) Nuisance Confounding.
When pieces of non-algorithmic context $\mathbf{U}_{t}$ do not meet all the components of the expertly targeted condition, they may not be non-algorithmic information at all – instead, they may be nuisance confounders. For instance, if $A_{t}$ is an advisor making a schedule recommendation, and $Y_{t+1}$ is next-term enrollment, then the $\textbf{U}_{t}$ of a student requesting schedule advice is a nuisance confounder, since students who request such advice are disproportionately likely to enroll the next semester regardless of the advisor’s actions. It is important to identify any nuisance confounding that might drive conditional dependence, by considering the implications of variable choice in the statistical test and looking for qualitative evidence of nuisance confounding. [^18] The distinction between a nuisance confounder and non-algorithmic information requires judgment, and a variable might be viewed as playing both roles at once; for instance, it arguably takes expertise for advisors to recognize a student is asking for help when those cues might not always be so obvious as an explicit request. Given the subjective nature of the distinction, we do not formally represent the distinction between nuisance confounders and non-algorithmic information in the model.
5 Quantitative Findings: Empirical Investigation of Advisor Expertise
We test for the conditional independence condition given in Proposition 4.2, which we showed is a necessary condition for expert targeting and sufficient for weakened components of it. Based on the results, we find five actions with significant conditional dependence on students’ outcomes. This test suggests the possibility of expertly targeted interventions, although we remind the reader of the impossibility result on identifying human expertise in targeting from quantitative observational data alone (see §4.3.1). Further, we build on the discussion in §4.3.4 and consider whether, in the MAAPS setting, alternative explanations (other than expertly targeted actions) may explain the conditional dependence results (§5.2.3), motivating the qualitative analysis in §6.
5.1 ExpertTest*: A Conditional Independence Test for Interventions
To test for advisor expertise in the MAAPS experiment, we adapt the test proposed by Alur et al. [5], which audits for expertise in algorithm-assisted prediction, to our setting of algorithm-assisted interventions. We refer to this test as the ExpertTest*. [^19] Just as for the theoretical framework, in our setting of algorithm-aided intervention, we replace the predictions of the original ExpertTest’s human forecaster with the advisor’s actions at a time $t$, $A_{t}$ and the predictive target $Y$ with their outcomes at time $Y_{t+1}$. We include the asterisk because of two differences in our setting: First, when predictions/actions may influence the outcome, the test is insufficient to demonstrate expertise, and second, because unlike Alur et al. [5], we also test for negative association for reasons described below.
Given a dataset $\mathcal{D}$ of $n$ $(\mathbf{x}_{t}^{i},a_{t}^{i},y_{t+1}^{i})$ observations, instantiations of the random variables ($\mathbf{X}_{t},A_{t},Y_{t+1}$), we look for independence of actions $A_{t}$ and outcomes $Y_{t+1}$ given the available data to an algorithmic predictor, $\mathbf{X}_{t}$:
$H_{0}:A_{t}\perp Y_{t+1}|\mathbf{X}_{t}$
If the ExpertTest* fails, then advisor actions are predictive of the student’s future outcomes in ways that cannot be explained by administrative data alone—suggesting they could be targeted to the student using non-algorithmic information. Indeed, the ExpertTest* corresponds to the statistical condition considered in §4.3 that is a necessary condition for expertly targeted action (Definition 4.1), and a sufficient condition for weakened components of it. Moreover, the test identifies potential expertly targeted actions relative to any algorithm based on $\mathbf{X}_{t}$, enabling its application even when risk scores are not recorded (as is the case for the MAAPS study).
5.1.1 ExpertTest* Implementation
For completeness, we briefly describe the implementation of ExpertTest* following Alur et al. [5], and note one point of departure from the original method.
We start by constructing a sequence of $K+1$ datasets that would be exchangeable under the null hypothesis $H_{0}$ given above, the first of which $\mathcal{D}_{0}=\{(x_{t}^{1},a_{t}^{1},y_{t+1}^{1})\ldots(x_{t}^{n},a_{t}^{n},y_{t+1}^{n})\}$ is the true dataset, and the other $K$ of which are synthetic datasets $\mathcal{D}_{1}\ldots\mathcal{D}_{K}$. To construct those synthetic datasets, we identify $L$ pairs of individuals with (nearly) identical $\mathbf{X}_{t}$ values (“algorithmically-indistinguishable individuals”) in $\mathcal{D}_{0}$. [^20] When individuals in $\mathcal{D}_{0}$ cannot be paired exactly, as is the case for our dataset, the test can be relaxed to accommodate approximate pairing, introducing error that scales in match quality (measured by a distance function on $\mathbf{X}_{t}$). We refer readers to [5] for the justification of the relaxation as well as rigorous statistical justification of the test. Then, using those $L$ pairs, $K$ synthetic datasets are constructed from $\mathcal{D}_{0}$ by making swaps of the actions advisors take for paired individuals with probability $\frac{1}{2}$.
The mean squared error of a dataset is defined as $MSE(\mathcal{D})=\sum_{i=1}^{n}(a_{t}^{i}-y_{t+1}^{i})^{2}$ with $n=|\mathcal{D}|$. Let $11[\alpha\gtrsim\beta]=$ 1 if $\alpha>\beta$ with ties broken randomly. Given exchangeability, the order statistics over the mean squared error $MSE(\mathcal{D})$ of these datasets,
$\tau_{K}=\frac{1}{K}\sum_{k=1}^{K}1[MSE(\mathcal{D}_{0})\gtrsim MSE(\mathcal{D}_{k})]$
will be distributed uniformly over $\{0,\frac{1}{K},\frac{2}{K}\ldots\}$. If the null is true, then, $\mathbb{P}(\tau_{K}\leq\frac{J}{K})\leq\frac{J+1}{K+1}$ for $J\leq K$. For $\tau_{K}=\frac{J}{K}$, then, we can reject the null with $p$-value $\frac{J+1}{K+1}$. Since $\frac{J+1}{K+1}\approx\tau_{K}$, we reject the null when $\tau_{K}$ is small.
Intuitively, the test assesses whether advisor actions predict outcomes in ways unexplained by algorithmically-visible features $\mathbf{X}_{t}$. Randomly swapping actions among individuals with identical $\mathbf{X}_{t}$ should not systematically worsen predictive error—as measured by $MSE(\mathcal{D})$— unless the original actions contained additional information about outcomes, whether through access to non-algorithmic information $U_{t}$, a causal effect of actions, or both.
As discussed in §4.3.3, actions intended to assist at-risk students may be negatively associated with outcomes. This is a key point of departure from the pure prediction setting [5], where good predictions must be positively associated with the outcomes. To address this, for every action $A^{i+}_{t}$ (e.g., recommending a schedule), ExpertTest* is also applied to the transformed action variable $A_{t}^{i-}=1-A_{t}^{i+}$ (e.g., not recommending a schedule). We define
$MSE^{-}(\mathcal{D})=\sum_{i=1}^{n}\left((1-a_{t}^{i})-y_{t+1}^{i}\right)^{2}$
and run the same procedure described above using $MSE^{-}$, allowing us to detect cases where taking an action is associated with worse outcomes, consistent with effective targeting of hidden risks. [^21] Another reasonable choice would be to run a two-tailed test on the order statistic $\tau_{K}$, which we opt against primarily for reasons of interpretability. See Appendix A.3.2 for a comparison; the methods yield similar results.
5.1.2 Application to MAAPS Data
The ExpertTest* is framed in terms of individual actions at a single timestep $t$ affecting outcomes at $t+1$, so to the multistage, multi-intervention nature of collegiate advising complicates variable definition. We run multiple different variants of the ExpertTest* to accommodate these complexities. Below we describe how we define $\mathbf{X}_{t}$, $A_{t}$, $Y_{t+1}$, and $n$ for these tests. All test parameters and variables, as well as details of how we chose test parameters $L$ and $K$, are summarized in Appendix A.3.3.
Algorithmic Information.
For all the tests, we take $\textbf{X}_{t}$, the algorithmic information, as students’ attributes at baseline excluding their race and ethnicity, since Georgia State University excludes this information from all its early warning systems [65]. Following our data cleaning process (§3.1.3), the population consists of 424 students having 3,507 meetings over four years (Figure 6).
The test requires matching students with approximately identical algorithmically-available features. We pair students greedily using a Euclidean distance function applied to fixed-range scaled features (Figure 7). [^22] Unlike the procedure for regressions (§7), we do not convert SAT scores to ACT scores, with the effect that students who took different college entry tests are not matched. This allows us to use students’ scores broken down by subject, for which there is no official conversion across tests. We select the number of pairs for each test, $L$, to yield an acceptable tradeoff between test power, which increases in $L$, and test error, which increases in match distance (Appendix A.3.3).
Actions.
We draw our action data from the lists of advisor interventions (described in §3.1.2).
To avoid an exponential blowup of actions, we binarize actions as any advisor applying intervention $i$ in a meeting with the student from the list of 20 over a particular timescale $t$. We denote this binary variable by $A_{t}^{i+}$. We pool intervention data from all advisors in order to capture potential use of external information by any advisor. Because of inconsistencies in the quality of Advisor 1’s intervention data (Appendix A.1.4), we re-run the test excluding students assigned this advisor and confirm the results’ consistency.
As discussed in §4.3.3, for actions to assist at-risk students where the risk factors are not encoded in $\mathbf{X}_{t}$, an action that supports students might be negatively associated with outcomes. Thus, we also run each test on $A_{t}^{i-}=|1-A_{t}^{i+}|$, which detects cases in which not taking the action predicts positive outcomes, or equivalently, taking the action predicts negative outcomes.
We test each intervention variable separately for simplicity, but in practice, interventions are often applied together by advisors. Their combined application may be important for their effectiveness or targeting. To account for this, we examine correlations between actions the ExpertTest* detects as significant, summarizing in Appendix A.3.6.
Outcomes.
Of the three outcomes we examine, persistence, graduation and GPA (at various timescales), all but GPA are naturally binary; we binarize the GPA outcome by testing for GPA above the mean of 3.5, testing more thresholds for robustness. [^23] We note that the MSE here differs from its normal interpretation—whereas MSE is normally a comparison of predictions of an outcome to that outcome, here we look at whether actions predict binary outcomes. Future work might explore alternate test statistics or conditional independence tests from Alur et al. [5].
Timescale.
The timescales of our action and outcome variables are chosen to avoid the introduction of feedback loops that could arise due to the multistage nature of the MAAPS experiment. [^24] For instance, if $A^{i}$ is having intervention $i$ performed at any point in 4 years, and $Y$ is persistence throughout those same 4 years, this introduces a backward edge $Y\rightarrow A^{i}$ because students who drop out will not meet with their advisors. We run tests over two different timescales:
  1. 1.
    Semesterly tests: Among enrolled students in each semester (Figure 6), we look at the result of the application of each intervention $i$ in a meeting that semester on term GPA or persistence to the next term (for spring terms, persistence is defined as enrollment/graduated status in either the summer or fall). In these tests, students’ semester of enrollment is added as a feature to their baseline attributes $\mathbf{X}_{t}$, and cross-semester matches of students are prohibited.
  2. 2.
    Early-intervention tests: Among students enrolled for the first year ($n=396$), we look at the result of each intervention $i$ that year on four-year graduation and four-year persistence.
Adjustment.
To account for multiple testing (40 hypotheses for each timescale and outcome, corresponding to $A_{t}^{i+}$ and $A_{t}^{i-}$ for $i\in[1,20]$), we adjust $p$-values for each timescale and outcome using the Benjamini-Hochberg procedure.
5.2 Results and Discussion
5.2.1 Results
For the ExpertTest* on semesterly timescales, four interventions demonstrate significant conditional dependence with GPA $\geq 3.5$ (Table 1) and one intervention demonstrates significant dependence with persistence (Table 2). Significance levels for all 20 interventions are given in Appendix A.3.5, alongside robustness checks demonstrating the consistency of the results to inclusion of potential transfer students, choice of GPA threshold, and exclusion of students assigned advisor 1.
These five significant interventions comprise 5,615 of the 8,385 total interventions taken by advisors following our data cleaning process. Accordingly, we estimate that 67%, or two-thirds, of interventions taken by MAAPS advisors are potentially expertly targeted. We provide this numerical estimate of the pervasiveness of significant interventions, with the caveat that a significant ExpertTest* result is necessary but not sufficient for expertly targeted action to be present. We address this limitation in §5.2.3 and §6, which provides additional qualitative evidence of expert targeting.
Two of these significance results have a positive direction, indicated by a (+)(+)( + ) in Tables 1 and 2. These correspond to tests on $A_{t}^{i+}$; the $p$-value for positive direction results should be interpreted as the probability that actions taken by advisor that cannot be explained by administrative data predict a positive outcome for the student (high GPA or next-term enrollment) just by chance. Three of the significant interventions have a negative direction, corresponding to tests on $A_{t}^{i-}$ and notated by ()(-)( - ), with $p$-values in these cases interpreted as the probability that actions taken by advisors that are unexplained by administrative data predict negative outcomes (low term GPA or failure to enroll next term) just by chance.
We note that the tests of first-year interventions on a four-year timescale do not meet the significance threshold (see results in Appendix A.3.5). We suspect this is due to the long time range of the hypothesis: Advisors’ actions in the first year are unlikely to contain detectable signal about outcomes three years later.
Semesterly tests on high GPA: top five interventions
Intervention Direction BH corrected $p$-value
20 (Other) (+) 0.039960*
16 (Schedule Appointment) (-) 0.039960*
2 (Correct Schedule) (-) 0.049950*
4 (Progression Issues) (-) 0.049950*
11 (Career Counseling) (+) 0.053280
Table 1: Top five most significant interventions for the ExpertTest* with $A^{i}_{t}$ indicating application of intervention $i$ during a given term, and $Y_{t+1}$ as that term’s GPA $\geq 3.5$ (the mean student GPA).
Semesterly tests on persistence: top five interventions
Intervention Direction BH corrected $p$-value
17 (Recommend Schedule) (+) 0.039960*
20 (Other) (+) 0.119880
16 (Schedule Appointment) (-) 0.333000
6 (See Instructor) (-) 0.399600
1 (Seek Academic Support) (-) 0.399600
Table 2: Top five most significant interventions for the ExpertTest* with $A^{i}_{t}$ indicating application of intervention $i$ during a given term, and $Y_{t+1}$ as persistence to the next term.
5.2.2 Discussion: Effect Sizes and Significance
Interestingly, there is a higher number of significant interventions for high term GPA than for persistence. We hypothesize three reasons for this trend. First, the power of the ExpertTest* is decreased for outcomes with high or low base rates, and persistence to the next semester is relatively common (Appendix A.3.4). Second, failure to persist is a less straightforward negative signal than a low GPA. For example, the comment data indicates a student might not re-enroll because they transferred to a school that better fits their needs, took a break to pursue a valuable opportunity outside of school, or faced family circumstances beyond their control. As a result, high-quality advisor support may not have a unilaterally positive effect on persistence. Third, the effects of actions advisors take to promote persistence might only surface over longer time ranges, whereas semester GPA can easily be impacted by advisors’ recommendations in a particular term (like recommendations to take particular courses).
Additionally, for all outcomes, it is possible that additional interventions meet the necessary condition for expertise but remain undetected. Just like outcomes, the ExpertTest* may not identify dependence for interventions with more extremal base rates—for instance, very rare interventions. This is because, for rare interventions, there are fewer “swaps” of algorithmically-indistinguishable individuals’ intervention values that can affect the mean squared error of a synthetic dataset, as most individuals’ intervention values are already equal to 0. Accordingly, Appendix A.3.4 shows that rarer interventions in the MAAPS dataset have fewer swaps between algorithmically-indistinguishable pairs that could change the MSE.
5.2.3 Discussion: Evidence of Expertise
We recall that the conditional dependence of actions and outcomes is only a necessary condition for expertise and that the test results do not confirm the five significant interventions are truly expertly targeted according to Definition 4.1. Expertly targeted action is in general not identifiable from observational data, as a consequence of Proposition 4.1. In other words, the conditional dependence of outcomes and actions can have alternative explanations than the use of non-algorithmic information, as our theoretical framework (§4.3) clarifies.
The set of interventions that are statistically significant raises three key questions to guide our subsequent qualitative analysis. Each of these questions corresponds to a particular concern about the mechanism that leads to conditional dependence, as discussed in §4.3.4: (A) non-algorithmic targeting, (B) supporting at-risk students and (C) nuisance confounding.
(A) Non-Algorithmic Targeting.
In the case of a significant result (actions predictive of outcomes in either direction), is there evidence that the interventions are non-algorithmically targeted? While many of the significant interventions present opportunities for advisors to incorporate non-algorithmic information into their decision to apply them, “Other” interventions (Intervention 20), which exhibit significant dependence with term GPA, make for an especially plausible instance of expert targeting. Since the list of interventions is already quite comprehensive, we hypothesize the decision to take “Other” actions might be tailored to bespoke student needs, identified through non-algorithmic information.
(B) Supporting At-Risk Students.
In the case the intervention is predictive of negative outcomes, is there evidence that advisors are targeting to at-risk students? Most of the significant interventions predict negative, rather than positive, outcomes. If advisors are not targeting on risk factors (and there is no other hidden confounding), then it is possible these actions of advisors hurt student outcomes. Since the interventions that negatively predict outcomes often signal that the student might be in trouble (like being advised about degree progression issues or advised to come in for an appointment), we hypothesize that advisors might be targeting based on algorithmically-invisible risk factors.
(C) Nuisance Confounding.
Could any significant results be driven by nuisance confounding? The association of Intervention 17 (being advised about recommended schedule) with persistence suggests some interventions might be confounded by nuisance factors. It may be that students who are likely to persist to next semester are also likely to ask their advisors for schedule advice in the first place, introducing dependence between schedule recommendations and outcomes that is not due to targeting.
We turn to qualitative analysis of the advisors’ logs of students interactions to answer these three questions.
6 Qualitative Analysis: Evidence for Non-Algorithmic Targeted Action
As described in §4 and §5, the quantitative observational data allows us to test for a necessary condition (conditional dependence) for expertly targeted actions, but does not allow us to distinguish expertly targeted actions from other mechanisms that may also be driving the conditional dependence. However, we have access to a rich dataset of qualitative data on advisors’ notes with their students that sheds light on these mechanisms.
We conduct a two-stage qualitative analysis to better understand advisor discretion and intervention practices, complementing our quantitative findings. The first stage involves open coding to identify themes across all advisor comments, to familiarize ourselves with the data and its limitations. The second stage focuses on a targeted analysis of comments linked to statistically significant interventions from the ExpertTest* to investigate potential evidence of expert targeting. In this section, we describe the methods for our qualitative analysis in §6.1 and how they complement our quantitative findings about advisor expertise in §6.2. In §7.1.2, we discuss results from the exploratory qualitative analysis in connection to our analysis of advisor heterogeneity.
6.1 Qualitative Methods
6.1.1 Exploratory Coding
In the first stage of our qualitative analysis, we conduct exploratory coding on all advisor meetings comments to understand the nature of advisor-student interactions as documented by those comments. Following removal of students due to our data cleaning (§3.1.3), we code a total of 1,594 comments left by advisors on meetings with students in the treatment arm. Using the Atlas.ti software, we apply both deductive codes based on our research questions and inductive codes that emerged from the data. To determine the codes, Authors 1, 2, and 6 first coded subsamples of comments independently, and then met to design a codebook, coordinating to add any additional codes as necessary while coding the entire dataset. We categorize codes into key thematic areas—academic, nonacademic, and relationship-building—further refining these with subcategories and attention to whether the advisor take a directive or exploratory approach.
We note that, while these codes provide valuable insight into advisor styles and the content of meetings, they are necessarily limited to what advisors chose to document, and may not fully capture the scope or nuance of each interaction. Details on our coding process are provided in Appendix A.5. Full code definitions and examples are provided in Appendix A.4.1.
The open coding process in the first round of analysis reveals key themes in advisor-student interactions. By examining the distribution of codes across advisors, student racial groups, and intervention types, we are able to explore variation in advising styles, assess consistency in how interventions were documented, and contextualize observed differences in advisor effectiveness and treatment heterogeneity (more in §7.1.2). This analysis also serves as a check on data quality, showing whether advisors’ written comments aligned with the interventions they reported taking.
6.1.2 Coding for Expert Targeting
We perform a systematic qualitative analysis of the advisors’ freeform comments to investigate the potential mechanisms behind significant result for the ExpertTest* (§5.2). We focus on comments of meetings where advisors reported taking at least one of the five statistically significant interventions. These codes of advisor-provided reasons enable us to assess whether the observed dependence between action and outcome plausibly reflect non-algorithmic information or these other factors.
Authors 1 and 2 used a spreadsheet to identify whether advisors explicitly mentioned taking the intervention and provided a reason for doing so (Table 15). For comments that met these requirements, these authors then independently coded the reasons advisors gave for comments associated with different interventions, before meeting to merge the codes and standardize their usage. We evaluate whether these codes indicate (A) the presence of non-algorithmic, outcome-relevant information guiding the choice of intervention, (B) evidence of targeting at-risk students, or (C) the presence of nuisance confounders (e.g., student requests), in accordance with the questions arising from our quantitative analysis. Full code definitions and examples are provided in Appendix A.4.1.
Two interventions require additional coding strategies. Recommending a schedule (Intervention 17) was taken as a default action for most students, which, as discussed in §4.1, trivializes the targeted action component of the expertise definition. Thus we additionally code for customization, identifying whether advisors tailored the content of their recommendations based on student-specific context. For "Other" interventions (Intervention 20), we develop a taxonomy of subtypes of interventions, reporting only those applied frequently enough to ensure consistent interpretation. Details of these specific coding strategies are included in Appendix A.5.1.
To reduce confirmation bias in identifying human expertise, we adopt a conservative approach: we interpret advisor reasoning when only clearly stated and apply the same coding procedures to a sample of statistically insignificant interventions for comparison. [^25] We select the two most insignificant interventions on the ExpertTest* for each outcome with greater than 100 instances of application; one of these overlaps for both outcomes, leading to three total interventions.
6.2 Qualitative Findings on Non-Algorithmic Targeting
We use the qualitative coding described above to answer three important questions about the mechanisms for a significant result in the ExpertTest*, as described in §5.2.3: (A) whether non-algorithmic targeting is present, (B) whether advisors support at-risk students in case of the intevention predicting negative outcomes, and (C) whether there is evidence of nuisance confounding. We include selected excerpts from advisors’ comments for illustration; comments are paraphrased as needed for clarity and to preserve anonymity. [^26] To protect student privacy, all gendered pronouns in comments have been changed to they/them.
(A) Advisors use diverse, detailed non-algorithmic context to target interventions.
Across all significant interventions, we see a consistent pattern: advisors incorporate non-algorithmic context into their decisions about when and how to intervene. For example, in Intervention 2 (advising a correction to a student’s schedule), advisors incorporate complex information, such as students’ course preferences and history, in their decision to suggest a schedule correction. In one case, an advisor recommends a student postpone a difficult course until the summer based on the student’s self-reflection and planning:
“Student stated that they are NOT feeling comfortable in their math course. This is a course that they need special attention on and they would like to take it in the summer. Because they have already taken two classes in the summer, I felt comfortable advising them to withdraw from the course.”
We note that advisors do not always rely on the student’s preference and, in some cases, use their discretion to provide recommendations against the student’s own preference:
“Student needs to add more credits to fall. Offered [MATH XXX]. Student didn’t want to take [MATH XXX] because they said they needed a ‘break’ from math. However, this will help the student stay on track, so I strongly encouraged them to take the class in the fall. Student says they will do so. ”
Unlike the first student, this second student needs to enroll in more courses and does not offer a clear reason for postponing the math course. In this case, the advisor encourages the student to take the class. Advisors, then, not only use non-algorithmic information to decide when to intervene, but also how to intervene, adjusting their particular recommendation—in this case, whether to advise postponing a course—based on context.
These customizations are especially crucial for schedule recommendations (Intervention 17), for which advisors consider detailed factors like student health conditions, motivation to stay in college, and plans to transfer in order to craft schedules for their students:
“Student has [a health] condition and has to make sure all of their courses are in proximity so as to avoid walking too much.”
“Student is in college because of pressures from their father, who could not attend college. Because they are covered by the GI Bill, the student did say they would give this first year their all. So my goal is to get them into classes next semester that will inspire them.”
“Student wants to transfer to GT [Georgia Tech]. Says they has conditional acceptance after 30 credit hours. I explained that not all CS courses transfer to GT and they need to find out what transfers. For spring, they want to take 18 credits to ensure they will reach 30 credit hours at the end of spring.”
While it is difficult to confirm these factors are not algorithmically predictable from basic administrative data or directly influence outcomes, the complexity of the reasoning behind deciding on an intervention and the variety of the reasons provided for those actions make this plausible in many cases.
The process of choosing how to act on non-algorithmic information also involves expertise. One piece of information could support multiple paths of action. The comments show how advisors exercise discretion to select an appropriate intervention from multiple options, based on the information they gather, attending to subtle details to decide the appropriate course of action. For instance, when a student is struggling with their coursework, advisors sometimes do not just recommend that students correct their schedule, as in the above cases, but instead warn the student about more dire progression issues (Intervention 4), based on their judgment of the severity of the problem. One advisor, for example, suggests a student reconsider their enrollment in the pre-nursing program based on their reported difficulty with their science courses:
“Student in pre-nursing major stated that math was giving them trouble. Without being discouraging, I explained the GPA requirements”.
Moreover, for any given intervention, the distribution of kinds of information that can persuade the advisor to act has a long tail: in most cases, there are many different kinds of information that can lead an advisor to take a particular action, making it difficult to improve a targeting algorithm by simple augmentations of the dataset and data collection process. Table 3 demonstrates the diversity of the reasons advisors cite for taking all the significant interventions we analyzed. When advising a schedule correction (Intervention 2), for instance, advisors consider matters including student’s course preferences, performance, financial aid requirements, projected course loads, prior placement exam performance, and credits earned elsewhere (which often have not been recorded in administrative data at the time of their conversation).
Finally, our analysis reveals that these patterns of plausibly non-algorithmic, diverse information supporting discretionary action are largely unique to the significant interventions (Appendix A.5.5).
(B) Evidence for targeted action supporting at-risk students.
Among the significant interventions that predict negative future outcomes, advisors take actions based on non-algorithmic information indicating potential risk factors to the student’s success.
Consider the intervention of recommending a student schedule an appointment (Intervention 16). Most often, advisors take this action for students who fail to schedule appointments or miss their scheduled appointments, an intuitive risk factor for poor outcomes:
“I called student because no appointment yet.”
“Student did not attend appointment. I called student to reschedule; student stated they are having to meet with financial aid due to their refund check sent to wrong bank.”
Advisors also take this intervention for students with exceptionally difficult personal or academic circumstances:
“They are currently living with friends and family. Apparently their mom is trying to figure out a new living situation and because of such, they do not have a permanent place to call home.”
“Student was dropped from courses back on Jan 18th due to non-payment. Student had not been attending courses the whole term. Student explained when we met the first time that they did not have the books.”
Table 3 shows additional plausible risk factors, like failure to enroll or poor course performance, guiding advisors’ requests to schedule an appointment and the other interventions that predict poor outcomes, recommending a schedule correction (Intervention 2), and warning about progression issues (Intervention 4). This supports the hypothesis that these actions negatively predict outcomes because advisors are performing these actions to support at-risk students that go unidentified by risk prediction algorithms based on limited administrative data, rather than that advisors are performing actions that have a negative effect on the students.
(C) Some interventions are confounded by nuisance factors.
While all significant interventions show signs of non-algorithmic targeting, some also show signs of a confounding effect, in which advisors perform interventions after students explicitly request they do. One such case is recommending a schedule (Intervention 17), which positively predicts next-term persistence ($p$=0.039), as students often seek scheduling advice for the next semester:
“NEUR student wanted to talk about spring semester and what to expect.”
“Student wanted to talk about college career.”
“Student came in the office to discuss schedule for SU ’17 and FA’17 semester.”
Such student requests confound the relationship between actions and outcomes with a factor that is arguably not non-algorithmic information. A student requesting schedule recommendations for next semester from their advisor likely also intends to enroll regardless of the advisor’s particular recommendations. Similar confounding also occurs in comments related to schedule corrections (Intervention 2), progression issues (Intervention 4), and being advised to schedule an appointment (Intervention 16), as the counts in row 12 of Table 3 (student-initiated interventions) demonstrate.
Significance on ExpertTest*: $\times$ $\times$ $\times$
Reason Category Int 2 Int 4 Int 5 Int 9 Int 10 Int 16 Category Sum
1 Course Credits and Placement 3 4 5 - - - 12
2 Course Problem 11 6 6 - - 6 29
3 Degree Progress - 8 3 - - - 11
4 Documents - - - 5 - - 5
5 Fin. Registration Problem - - - 52 - 4 56
6 Load Management 18 - 1 - - - 19
7 Major Choice - 8 1 - - 18 27
8 Non-fin. Registration Problem - - - - 3 - 3
9 Personal Circumstances - 2 - 1 - 5 8
10 Low Engagement - - - - - 109 109
11 Scholarship Process 2 - - 6 - 1 9
12 Student-Initiated 3 - - 1 - 8 12
13 Transfer/Dropout - 3 - - - 3 6
14 Unregistered - - - - - 12 12
Table 3: Counts of specified reason categories for each intervention type mentioned during advising meetings, indicating how frequently each reason was associated with a particular intervention, for Intervention 2 (Correct Schedule), Intervention 4 (Progression Issues), Intervention 5 (Adjust Course Balance), Intervention 9 (Address Financial Issues), Intervention 10 (Address Non-Financial Collegiate Requirements), and Intervention 16 (Schedule Appointment).
7 Heterogeneity in Student Outcomes
Beyond studying expertly targeted actions and its mechanisms, in this section, we explore the broader question of impact: In what ways does discretionary decision making affect student outcomes and equity? Motivated by recent work on the equity implications of algorithmic interventions in education [23, 24], we focus our analysis on how treatment effects (on graduation outcomes) vary across subgroups defined by (1) the primary advisor assigned to students; and (2) student demographics, specifically race and ethnicity. To estimate these heterogeneous treatment effects, we fit logistic regression models, as well as causal forest models to account for potential non-linear interactions. All models are trained on data from 745 students (392 in the treatment group and 353 in control). [^27] For these regressions, we exclude students identified as multiracial and those categorized as non-resident aliens, as these groups are substantially smaller in size compared to other demographic categories (see Appendix A.6.1 for a complete list of exclusions).
Across advisors, we observe some heterogeneity in graduation outcome effects (see Table 18), though the differences are not statistically significant. However, a qualitative analysis of advisor comments in §7.1.2 suggests variations in advising approaches. Synthesizing both analyses, we find preliminary evidence that student outcome variations may be associated with advising style: students assigned to advisors with a holistic approach tended to fare better than those with academically-focused advisors.
The original MAAPS study (§3.1.1) noted potentially greater treatment effects for Black students. [^28] After six academic years, Black students experienced a more positive treatment effect on graduation outcomes from the MAAPS intervention than non-Black students [65]. Building on this, we conduct a more systematic analysis of heterogeneous treatment effects across all race and ethnicity groups. We fail to find significant heterogeneous effects by race-ethnicity after correcting for multiple hypothesis testing (see results in Table 20). We remark, however, that the original MAAPS study is not well powered to detect small differences in treatment effect (see discussion in Appendix A.6.5), and the lack of statistical significance need not imply the absence of heterogeneous treatment effects.
7.1 Heterogeneous Effects Across Advisors
We look for heterogeneous effects across advisors.
7.1.1 Quantitative Analysis
Each student in the MAAPS study treatment group was assigned to one of the five different advisors. For each of the advisors $j\in\{1,\ldots,5\}$, let $A_{ij}$ denote whether student $i$ was assigned to advisor $j$ [^29] Some students had interactions with multiple MAAPS advisors. For these students, we let $A_{ij}=1$ if most of student $i$’s interactions are with advisor $j$ and we set $A_{ij^{\prime}}=0$ for all $j^{\prime}\neq j$., $Z_{i}\in\{0,1\}$ denote whether the student has successfully graduated, and $T_{i}\in\{0,1\}$ indicate whether the student was assigned to the treatment group. We fit the following logistic regression to model the graduation outcome, [^30] We additionally considered exploratory analyses for the cumulative GPA outcomes. However, due to substantial selection bias in those models (only enrolled students have a GPA), regression results may not provide reliable or generalizable inferences. controlling for the baseline student attributes (gender, high school performance reflected in the composite ACT, financial status in terms of expected financial contribution, and academic preparation measured by other credit hours, which we describe further in Appendix A.1.1):
$\begin{split}\text{logit}&\big{(}\mathbb{P}(Z_{i}=1)\big{)}=\,\beta_{0}+\beta_{1}T_{i}+\sum_{j}\theta_{j}A_{ij}+\gamma_{1}\text{(Gender)}_{i}\\ &+\gamma_{2}\text{(ACT)}_{i}+\gamma_{3}\text{(EFC)}_{i}+\gamma_{4}\text{(Hours\_Other)}_{i}\end{split}$ (1)
We do not find significantly non-zero heterogeneous treatment effects across advisors (see Appendix Table 18). However, both logistic regression and causal forest models produce a consistent ranking of advisor-specific effects (Appendix Table 19). In the logistic regression model, Advisor 3 has the highest advisor-specific coefficient $\theta_{3}=0.51$, followed by Advisors 2 and 1 ($\theta_{3}$ of 0.02 and $-0.25$ respectively). Similarly, the causal forest analysis estimates the highest heterogeneous treatment effect for Advisor 3, with an effect size of 0.11. The ensuing qualitative analysis sheds light on potential connections to differences in advising style.
7.1.2 Qualitative Analysis
Figure 2: Normalized distribution of qualitative codes (meeting topics) for each MAAPS advisor. Advisors 1 and 2 focus on academic topics, while Advisor 3 emphasizes relationship-building and nonacademic discussions during student meetings. The Generic Suggestion code here refers to the advisor asking the student to follow up with them later without specifying a reason (see full details in Appendix A.4.1).
Despite the lack of significant effects in the quantitative analysis, the qualitative analysis of the advisors’ comments (§6.1.1) reveals notable differences in their approaches to advising. In particular, we highlight two distinct advising styles that primarily differ in the scope of topics discussed in advisor-student interactions. [^31] We note that our analysis is limited to advisors’ written comments, which might not reflect the complete range of applied advising strategies. The first style (reflected in the comment of Advisors 1 and 2) is academic-oriented, in which the advisor adheres to a focused set of actions and topics in meetings including academic performance, degree progression and requirements. On the other hand, the second style (Advisor 3) is holistic, with broader engagement in non-academic matters and student’s interests and emotional states. [^32] Advisors 4 and 5 are excluded from our analysis due to the limited number of comments, which did not provide enough data for a meaningful comparison. Figure 2 shows the full distribution of subject matter codes for each MAAPS advisor.
To illustrate the academic-focused style of advising, we turn to Advisor 1, whose comments consistently center on academic planning, frequently including references to degree audits, major maps, and specific course planning. Advisor 1’s comments follow a structured pattern across most student meetings, reflecting a more routine approach to academic advising: [^33] Advisor 1 utilized a standardized comment template in 76% of all meetings (647 out of 849 total comments). Of these repeated comments, 440 include ”Additional comments”, although only 44 specifically addressed unique student concerns.
"Student came to the office to discuss schedule for SP’18 semester, printed and reviewed degree audit, student was shown major map, reviewed EAB for success markers, and notified student…"
Interestingly, we notice that Advisor 1 is the only advisor who explicitly refers to the use of algorithmic tools (“reviewed EAB for success markers”) during advisement.
When Advisors 1 and 2 do personalize their advice to the student, their scope of consideration is also largely constrained to the student’s academic status. For example, Advisor 2 writes:
"Student wants to change major from bio to exercise science… I advised that they could possibly do mini-mester for spring in order to compensate for credits lost towards degree."
Advisor 3 exemplifies a more holistic advising style. The comments left by Advisor 3 reflect a broader engagement with the student as a whole person, addressing not only academic planning, but also emotional well-being, personal motivations, and long-term goals. The advisor is also more likely to note and adjust their advice to the specific concerns and needs that students bring up during each meeting, even if these lie outside the scope of academics.
In one example, Advisor 3 encourages a student to reflect more deeply on their desire to help others, guiding them to explore opportunities that align with their interests:
"The student wants to help people, but I challenged them to think about why and how. I advised them to look into some organization fairs where they can start getting experience working with people."
Advisor 3’s comments also reveal a pattern of empathic support and thoughtful challenge, which is also reflected in a significantly higher proportion of relationship-building qualitative codes applied to each student meeting (see Figure 2). For example, Advisor 3 responds to a student experiencing anxiety about their future by acknowledging their concerns, exploring possible academic paths, and pointing them toward mental health resources:
“Student is very anxious about GSU, college and career in general… I challenged them to get exposure to the field. They did ask about the counseling center, for which I provided them with resources.”
The structure and content of these conversations vary considerably by advisor. As shown in Table 4, Advisors 1 and 2 have relatively high average Jaccard Similarity Index (JSI) scores,343434 For a pair of meetings, the Jaccard Similarity Index (JSI) is defined as the ratio of the intersection to the union of the codes applied to each meeting. For each advisor, we calculate the average JSI for all pairs of their meetings to assess how similar their comments were overall, with a higher average indicating less diversity in comments. reflecting a repetitive use of qualitative codes across sessions, consistent with a structured advising approach. In contrast, Advisor 3 has a notably lower JSI value, indicating diversity in the topics and strategies employed, which aligns with a personalized advising style.
Advisor # Comments Average JSI
Advisor 1 843 0.3006
Advisor 2 364 0.3791
Advisor 3 395 0.1914
Table 4: Average pairwise Jaccard Similarity Index (JSI) on qualitative codes applied in the meetings for each advisor.
The quantitative analysis of heterogenous treatment effects suggests that Advisor 3 has a more positive impact on student outcomes than Adviors 1 and 2. This observation may reflect the advantages of a more holistic advising style compared to the academic-oriented approach for this particular college advising context. In our setting, it is also important to note the limited power to detect heterogeneous effects and the potential for unmeasured effects, such as student satisfaction or self-efficacy, which may still be influenced by the advising styles but are not captured in the graduation-based quantitative analysis. Future work is required to further investigate whether certain advising styles lead to more student success.
7.2 Heterogeneous Effects by Student Demographics
To examine heterogeneity in treatment effects across student subgroups, we estimate a logistic regression model predicting the binary graduation outcome $Z_{i}\in\{0,1\}$, where $Z_{i}=1$ indicates that student $i$ graduated. The main treatment indicator $T_{i}\in\{0,1\}$ denotes assignment to the treatment group. For the independent variables, we include student race-ethnicity indicators (Black, Asian, Hispanic [^35] The indicator variable White was omitted from the model to avoid multicollinearity.), their interactions with treatment, and baseline covariates (as controls), using the same specification as Model (1).
$\begin{split}\text{logit}\big{(}\mathbb{P}(Z_{i}=1)\big{)}=\,&\beta_{0}+\beta_{1}T_{i}+\sum_{r\in\mathcal{R}}\left(\theta_{r}\cdot\text{Race}_{r,i}+\delta_{r}\cdot T_{i}\cdot\text{Race}_{r,i}\right)\\ \quad+\gamma_{1}\text{(Gender)}_{i}+&\gamma_{2}\text{(ACT)}_{i}+\gamma_{3}\text{(EFC)}_{i}+\gamma_{4}\text{(Hours\_Other)}_{i}+\epsilon_{i}\end{split}$ (2)
We focus specifically on subgroup effects by race, motivated by Georgia State University’s institutional commitment to improving educational access and equity for historically underrepresented students [61], as well as the broader concerns that early alert and predictive analytics systems can unintentionally hurt educational outcomes for minority students [23]. We note that race is not used as an input in the predictive analytics system deployed at GSU.
We do not find statistically significant interaction terms between treatment and race variables, after correcting for multiple hypothesis testing (see results in Table 20). This may not be an indication of the lack of heterogeneous treatment effects, as the MAAPS experiment may not be well-powered to detect small heterogeneous treatment effects. We run a power analysis in Appendix A.6.5 and find that, assuming a desired power of 80% and a sample size of 745 students from our experiments, the minimal additive detectable effect is 18%, meaning that our study on the MAAPS dataset can only detect a large increase in the graduation rate. Furthermore, from Table 17, we can estimate that in order to achieve significant increase of the average graduation of $\sim$ 5% within the subgroup of Black students, we would require a total population sample size of approximately 10,000 students.
Additionally, to assess whether advisors’ intervention choices were systematically influenced by students’ race and ethnicity, we conduct a series of logistic regression analyses. For each of the five significant interventions (listed in §5.2.1), we estimate two models (see Appendix A.6.8): the first that included race-ethnicity variable alongside all other baseline student characteristics (listed in Appendix A.1.1), and the second model that excludes this variable. Across all cases, we find no difference in predictive performance between these two classifiers (see Table 22), suggesting that, conditional on observed student characteristics, advisor decisions to select particular interventions during meetings were not meaningfully shaped by student race and ethnicity.
Together, these findings highlight the need for further investigation into how discretionary decision making interacts with equity in student support.
8 Discussion
Our study of the MAAPS intervention shines a light on the central role of discretionary action by advisors in the use of an algorithmic decision support for advising. This role for discretion raises key questions for future algorithmic systems, introduces new challenges for existing methodology, and suggests the importance of reflecting on the normative role of discretion in institutional decision making.
8.1 Future Algorithmic Decision Systems: Design, Implementation, and Evaluation
The central role of discretionary decision making in the MAAPS study has important implications for the design, implementation, and evaluation of algorithmic decision aids, in advising and beyond.
8.1.1 Designing for Human Discretion
Many examples of non-algorithmic information used by MAAPS advisors point to sources of data that are inherently difficult, if not impossible, to encode quantitatively. Advisors, for instance, frequently reference psychological constructs such as student engagement and motivation to justify their interventions, which are notoriously hard to measure [25, 26], but can be observed by advisors through conversation. Moreover, consistent with findings from other domains where human discretion helps compensate for flawed or incomplete data [17], advisors in the MAAPS experiment also draw on non-algorithmic information to fill gaps in administrative data that risk prediction tools use—such as missing course credit information or delays in placement test results.
Even if technologies like language models can gather some of this information from students, interpersonal relationships may still be essential for eliciting sensitive information and for providing meaningful support [60]. Moreover, expanding structured data collection about students also raises privacy concerns that discretionary discussion with advisors can avoid.
Just as there are obstacles to encoding this information in an algorithmic system, these systems may also struggle to learn how to use this kind of information appropriately. Advisors often construct elaborate causal narratives about a student’s circumstances and likely behavior to guide their interventions. Even routine suggestions—such as advising a student to drop a course—can require intricate causal reasoning: how past coursework may affect future scheduling, why a student wants to withdraw, and, if the student plans to retake the course, whether their obstacles to succeeding will apply in the future. [^36] For example: “The student wanted to withdraw from their economics class, as their first exam was a 40%. They did admit they just didn’t study, and they understand the importance now. Because they brought in AP credit, they can withdraw and still be on track. But I made it clear that they cannot withdraw moving forward or else they’ll be forced to take summer classes to stay on track. They understand.” While human advisors naturally draw on this kind of causal reasoning—as has been observed across decision-making contexts [40]—current algorithmic systems are not built to learn these causal relationships [60]. [^37] Such as the Navigate tool [23] or the student-support chatbots used at GSU [52]. Moreover, our qualitative analysis of non-algorithmic information also underscores the importance of discretion in handling rare cases, a task at which even state-of-the-art language models often struggle, both in general and in advising contexts specifically [39, 44]. Rather than aiming to replicate human discretion, future systems might be more effective if designed to support and enhance human expertise [48]. This could be done by training new advisors to develop expertise by embedding expert strategies in large language models [76], or by automating routine administrative actions that do not require discretionary expertise (discovered through techniques such as those in the current work), thereby allowing advisors to focus on building relationships with students and exercising judgment where needed.
8.1.2 Implementation Challenges of Generalizing and Scaling
The MAAPS study’s positive treatment effects for graduation and credit accumulation at GSU shows the promise of algorithmic decision aids to improve educational outcomes. However, as with many educational interventions, MAAPS faces two key challenges to external validity: generalizing the intervention to new populations and scaling it across institutional contexts [42, 72, 37, 46]. These challenges are exacerbated in studies of algorithmic decision aids where human decision making plays a central role, like the MAAPS study. In such cases, positive outcomes cannot be neatly attributed to the decision aid alone; they also depend on advisors’ ability to skillfully mediate between the system and the students. Thus, scaling this intervention requires not only scaling the algorithmic component but also hiring similarly skilled human advisors who can use the tool. These discretionary interactions are difficult to replicate in new populations or standardize across institutional contexts, complicating both generalization and scalability.
Generalizing the intervention is complicated by both institutional culture and student demographics. At GSU, the MAAPS protocol was a natural extension of the schools’ existing advising approach, which likely facilitated its adoption by advisors. By contrast, advisors at other institutions may have been less receptive to the protocol, particularly if it departs from their advising norms [65]. Moreover, at GSU, a plurality of students are African American [19], which may shape how students experience predictive interventions. The risk of racialized stigma to being flagged as high-risk by an algorithm [75] may increase significantly in contexts where structurally disadvantaged groups are numerical minorities, especially if demographic characteristics are included as inputs in the predictive model (as they are not at GSU) [23].
Similarly, implementation challenges can blunt the effectiveness of the MAAPS intervention when scaled. Institutions that do not fully adopt the MAAPS advising protocol are unlikely to realize the same benefits from algorithmic decision aids as those observed at GSU. The MAAPS study itself illustrates this point: other participating institutions in the study struggled to implement core components of the intervention, such as assigning students to primary advisors from a centralized advising office rather than across academic departments [2]. None of these institutions achieved the same positive treatment effects as GSU, which the study authors partly attribute to these implementation challenges [65].
Another likely contributor to the success of the MAAPS intervention, and to predictive analytics more broadly at GSU, is the university’s investment in increasing the capacity of the human decision-makers. Prior work has shown that increased capacity in human outreach programs leads to larger effect sizes and higher power in RCTs [13], supported by the results of the RCT at GSU versus the other participating institutions. As part of its advising strategy, GSU reduced its student-to-advisor ratio from 700:1 to 300:1 [63]. Within the MAAPS intervention treatment arm, this ratio dropped to approximately 150:1 [65]. Many of the other institutions participating in the MAAPS were unable to consistently maintain this ratio due to advisor retention problems [65]. These numbers suggest that the effectiveness of decision aids may depend on the institutional ability to support and retain human advisors. Schools that adopt algorithmic systems as cost-saving measures—placing human decision-makers in resource-constrained situations or pressuring them to follow algorithmic recommendations [40]—are unlikely to replicate positive outcomes.
8.1.3 Evaluating Algorithms: Human Decision-makers as a Study Condition
The importance of human discretion to advising also has implications for designing future studies of decision support systems. Without understanding how the interaction between humans and algorithmic tools drives positive effects, replicating or scaling those effects of an implementation is difficult. Indeed, randomized field trials may depend on factors about the human’s cognitive response that could change outside the experimental setting—for instance, an increase in the false positive rate of algorithmic decision aid might influence how much human decisionmakers rely on it [59]. Thus, as part of developing randomized trials of real-world algorithm deployments [36], we argue it is important to develop techniques that build a thorough understanding of how the human decision-maker reacts to the decision aid.
Some researchers have attempted to study human discretion through controlled experiments with crowdworkers [29, 16, 58, 30], but such studies often lack external validity, failing to capture the knowledge of real decisionmakers and the complexity of institutional contexts [50]. By contrast, randomized field experiments can look at the kinds of discretion that arise in the wild, using a mix of quantitative and qualitative methods. Randomly assigning advisors to students in treatment, for instance, could help disentangle the effects of human discretion from those of the algorithmic tool itself. Additionally, studies of algorithmic decision aids should collect data—both structured and unstructured—on how humans interpret and act on algorithmic recommendations. The MAAPS study offers a promising model in its documentation of advisor actions and rich comments on meetings; this detailed picture of advisor decision making allows us to take a uniquely granular look at factors influencing advisor decision making in this paper. This data could be further improved by collecting records of the algorithmic outputs themselves and descriptions of how advisors used them.
8.2 Methodological and Practical Limitations in Capturing Human Expertise
Our study of discretionary decision making within the MAAPS dataset has two key limitations, both of which highlight broader challenges for future research on discretionary algorithm-supported decision making. First, our theoretical framework is designed for single-stage decision making, but collegiate advising (as in the MAAPS experiment) is a multistage process. Advisors work with students across multiple advising meetings, helping them achieve long-term academic and personal goals. As a result, the definition of expertise presented here (and the statistical condition used for the ExpertTest*) does not capture forms of human expertise that may not be immediately observable at one stage of decision making. For instance, an expanded notion of expertise might capture trajectories of actions leading to good outcomes. This raises a broader challenge in understanding discretionary decision making when the effects of decisions may be cumulative or take a long time to materialize.
Second, our study is limited by the lack of access to algorithmic outputs, such as the risk scores used in the MAAPS intervention. Direct access would enable us to understand advisor expertise relative to actual risk assessments. While our use of the ExpertTest* allows us to assess whether advisors target interventions for any hypothetical algorithm, it constrains the types of analyses we can perform. For instance, we did not incorporate recurring time-series data (Appendix A.1.2), as the test is not designed for such data. Future research could build on this work by collecting algorithmic outputs and exploring alternative dependence tests.
Finally, nuisance confounding, such as explicit student requests influencing advisor actions, hinders the ability of our statistical test to indicate potential expertly targeted action. This limitation underscores the inherent hazards of using observational data for studying causal questions and the importance of creating evaluation techniques that can build a causal understanding of discretion in decision making through controlled experimentation.
8.3 Functional and Normative Roles of Human Discretion in Institutional Decision Making
Our study provides suggestive (though not confirmatory) evidence that students achieve better educational outcomes due to advisors’ discretionary contributions. This finding contributes to a long-standing debate across decision-making domains about whether human discretion helps or hinders the delivery of high-quality interventions, and when its institutional role should be increased, maintained or diminished in light of automated decision systems.
Evidence across domains suggests that discretion is a double-edged sword. On the positive side, discretion enables flexible, “street-level” judgments by human decision-makers, allowing them to override algorithmic recommendations based on rich, real-time information to meet evolving goals [4, 34, 40, 68]. Discretion could also promote equity; for example, in child maltreatment call screening, call workers have helped close racial disparities by adjusting how they respond to algorithmic risk scores, mindful that the data driving these tools may reflect systemic biases [17, 27]. In addition to these benefits, our study suggests that advisor discretion acts as a bridge between predictions (e.g., risk scores) and effective interventions [47]. By eliciting and responding to non-algorithmic information from students, advisors select and deliver meaningful interventions that go beyond the capabilities of an algorithm. Whether advisor discretion also provides additional equity benefits—such as counterbalancing systemic biases of risk scores—remains an important question for future research.
On the negative side, discretion can hinder decision making, particularly in settings where decision quality is thought to depend mostly on prediction accuracy [41] (though this is not our setting). In some predictive contexts, algorithms may outperform human forecasters [20] and so lead to improvements in decision quality. Even when human decision-makers have access to information beyond what the algorithm does [5], studies show that people often misjudge when to override or defer to algorithmic advice, leading to over-reliance or under-reliance [9]. Moreover, discretion can introduce bias, as demonstrated by the case of a pretrial risk assessment tool in Kentucky, where judges were more likely to override low-risk scores for Black defendants than for white defendants [3]. Our finding that the MAAPS study is underpowered to detect heterogeneous effects for minority students is thus a notable limitation, as it constrains our ability to assess whether discretion in advising contributed to student outcome disparities in unintended ways.
Importantly, questions about discretion are not only about effectiveness but also about institutional values. The use of proprietary, opaque algorithms without a human in the loop can create accountability gaps, limiting the ability of students or advisors to contest decisions [15]. In other domains, it has been shown that rigid institutional mandates to follow algorithmic recommendations can undermine worker autonomy, generate resentment, and shift the balance of power away from human workers [68, 14]. Finally, those classified as “high-risk” by risk prediction tools may face stigmatization [75], and may express strong preferences for discretion over algorithmic decision making [38]. Particularly in educational contexts like advising, where personal connection and trust are critical, designing algorithmic systems must involve careful reflection on when and how to preserve human discretion.
Acknowledgments.
We would like to thank the hard work of the entire MAAPS project team at GSU. The efforts of the MAAPS advisors, particularly the diligence and dedication of Emily Buis, were instrumental in developing a rich source of data that made this work possible. We thank Timothy Renick for his visionary leadership at NISS and for sharing insights that provided valuable context and inspiration for this research collaboration. We also thank Tim Fulton, a prior staff member at NISS, for his support, as well as Andy Zhang for his research assistantship in the initial stages of this study. We thank Manish Raghavan for valuable discussions on his earlier work.
Research reported in this manuscript was supported by an Amazon Research Award Fall 2023 (L.T.L. and K.S.). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of Amazon.
Appendix A Appendix
A.1 MAAPS experiment and data
A.1.1 Baseline Characteristics
For each student, the dataset provides values of the following variables at time of study enrollment:
  • Student ID
  • Month and year of birth
  • Gender (selected from one of the following options: female, male, other)
  • Race/ethnicity (selected from one of the following options: non-resident alien, Hispanic of any race, American Indian or Alaska Native, Asian, Black or African-American, Native Hawaiian or Other Pacific Islander, White, Two or More Races)
  • Whether the student is a transfer student
  • Whether the student received a high school diploma or equivalent
  • The student’s high school GPA
  • The student’s SAT score, both composite and broken down by verbal and math sections
  • The student’s ACT score, both composite and on the English and Math sections
  • Number of college credit hours already earned by the student (for instance through AP classes)
  • The student’s expected financial contribution for the 2016-2017 academic year (in twelve \$$\$1050$ intervals, ranging from \$$\$0-\$1050$ to \$$\$22,051$ or more)
A.1.2 Timeseries Data
For each term, the dataset provides the value of the following variables for all students:
  • Gender
  • Graduation term and year (if relevant)
  • Term enrollment status
  • Term GPA
  • Cumulative GPA
  • Declared majors and minors
  • Hours of credit attempted and earned in the term, broken down by remedial and non-remedial credits
  • Total number of non-remedial credits earned up to and including the term
  • Student residency in university-owned housing during the term
  • Student participation in another intensive advising program
  • Student residency status for tuition purposes
  • Student’s assigned MAAPS advisor
A.1.3 Meeting Logs
For students in the treatment arm, MAAPS advisors log every student interaction with the following information:
  • What initiated the advising interaction with the student (one option selected from: registration or program review, advising system alert, regularly scheduled advising, early alert or instructor initiated, and student intiated)
  • Method of contact (one option selected from: in person, email, telephone, group, online, text, student’s primary/departmental advisor, other)
  • Intervention(s) performed in the meeting (any number selected from a list of twenty, described below)
  • Freeform comments on the meeting
The list of twenty interventions included:
  1. 1.
    Advised to seek academic support
  2. 2.
    Advised to correct schedule
  3. 3.
    Advised to declare or change major
  4. 4.
    Advised about progression issues
  5. 5.
    Advised to adjust course balance
  6. 6.
    Advised to see instructor
  7. 7.
    Advised to seek academic coaching
  8. 8.
    Advised to seek departmental or faculty guidance
  9. 9.
    Advised to address financial issues
  10. 10.
    Advised to address non-financial collegiate requirements
  11. 11.
    Advised to seek career counseling and service
  12. 12.
    Advised to seek engagement opportunities on campus
  13. 13.
    Advised to see professional counselor
  14. 14.
    Advised to seek disability services on campus
  15. 15.
    Advised to see medical professional
  16. 16.
    Advised to schedule an appointment
  17. 17.
    Advised about recommended schedule
  18. 18.
    Advised to see primary program advisor
  19. 19.
    Advised through the in-person development or review of the academic map/degree planner
  20. 20.
    Other (Describe in comments)
A.1.4 Data Quality
Figure 3: Normalized distributions of applied interventions per advisor, demonstrating an especially low diversity of interventions applied by Advisors 1 and 2. Advisors 4 and 5 are excluded due to the limited number of available comments for each.
There was some inconsistency in the way the advisors documented the interventions they performed during these meetings. In some cases, while there is clear evidence in the free-form comments that advisors performed a certain intervention during an interaction, they failed to mark the corresponding intervention boxes on the provided checklist. Advisor 1 in particular consistently selected limited sets of interventions (only ever indicating they had applied 10 of the 20 total interventions and only indicating they had applied four of them more than 100 times), even when their comments indicated that they had applied additional interventions. We show how the distribution of interventions applied varies for each advisor in Figure 3.
A.2 Theoretical Framework
A.2.1 Proof of Proposition 4.1
Proof.
We prove Proposition 4.1 by defining two SCMs $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ such that both have the same joint distribution over $(A_{t},Y_{t+1})$ but only $\mathcal{M}_{1}$ meets the definition for expert targeting. Without loss of generality, we construct $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ such that there is no algorithmically available context $X_{t}$, corresponding to a setting where the algorithmic recommendations are not followed by the human decision-maker. [^38] We note the example can be modified to include $X_{t}$, for example by distributing it as a Bernoulli and adding a term to the equations for $Y_{t+1}$ and $A_{t}$ multiplying the entire expression by $X_{t}$ in both models.
Let $\textsf{Ber}(p)$ denote a Bernoulli random variable with mean $p$ and let $p\in(0,1)$. Consider the following two SCMs:
$\mathcal{M}_{1}$
$\epsilon_{A_{t}}$ $\sim\textsf{Ber}(p)$
$\epsilon_{U_{t}}$ $\sim\textsf{Ber}(p)$
$\epsilon_{Y_{t+1}}$ $\sim\textsf{Ber}(p)$
$U_{t}$ $:=\epsilon_{U_{t}}$
$A_{t}$ $:=\max(U_{t},\epsilon_{A})$
$Y_{t+1}$ $:=\min(A_{t},\max(U_{t},\epsilon_{Y_{t+1}}))$
$\mathcal{M}_{2}$
$\epsilon_{A_{t}}$ $\sim\textsf{Ber}(2p-p^{2})$
$\epsilon_{Y_{t+1}}$ $\sim\textsf{Ber}(\frac{p^{2}-p-1}{p-2})$
$A_{t}$ $:=\epsilon_{A}$
$Y_{t+1}$ $:=\min(\epsilon_{Y_{t+1}},A_{t})$
First, we show that $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are observationally equivalent. Since the SCM’s probability distributions are defined over the noise variables $\{\epsilon_{V}:V\in\mathbf{V}\}$, Table 5 demonstrates how the exogenous noise settings lead to different settings over model variables $\mathbf{V}$.
To derive the observational distribution $\mathbb{P}(A_{t},Y_{t+1})$ for each model, we sum the probabilities of all exogenous variable configurations that result in each setting of $A_{t}$ and $Y_{t+1}$. Table 6 presents the distribution for both models. Since the results are identical, this shows the models are observationally equivalent with respect to $\mathbb{P}(A_{t},Y_{t+1})$.
Noise table for $\mathcal{M}_{1}$
$\epsilon_{A_{t}}$ $\epsilon_{U_{t}}$ $\epsilon_{Y_{t+1}}$ $U_{t}$ $A_{t}$ $Y_{t+1}$ Probability
0 0 0 0 0 0 $-p^{3}+3p^{2}-3p+1$
0 0 1 0 0 0 $p^{3}-2p^{2}+p$
0 1 0 1 1 1 $p^{3}-2p^{2}+p$
0 1 1 1 1 1 $p^{2}-p^{3}$
1 0 0 0 1 0 $p^{3}-2p^{2}+p$
1 0 1 0 1 1 $p^{2}-p^{3}$
1 1 0 1 1 1 $p^{2}-p^{3}$
1 1 1 1 1 1 $p^{3}$
Noise table for $\mathcal{M}_{2}$
$\epsilon_{A_{t}}$ $\epsilon_{Y_{t+1}}$ $A_{t}$ $Y_{t+1}$ Probability
0 0 0 0 $\frac{-p^{4}+4p^{3}-6p^{2}+4p-1}{p-2}$
0 1 0 0 $\frac{p^{4}-3p^{3}+2p^{2}+p-1}{p-2}$
1 0 1 0 $p^{3}-2p^{2}+1$
1 1 1 1 $-p^{3}+p^{2}+p$
Table 5: Noise tables for models $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. These tables show the values of the endogenous variables under different exogenous noise settings and their associated probabilities.
$A_{t}$ $Y_{t+1}$ $\mathbb{P}(A_{t},Y_{t+1})$
0 0 $p^{2}-2p+1$
0 1 00
1 0 $p^{3}-2p^{2}+p$
1 1 $-p^{3}+p^{2}+p$
Table 6: Observational distribution over $A_{t}$ and $Y_{t+1}$ in both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, obtained by summing over the rows of the noise table that produce these settings of $A_{t}$ and $Y_{t+1}$ for each model.
Next, we show both of these models meet our assumption that SCMs are causally minimal. Causal minimality is equivalent to that, for all causal parents of $Y$, $X\in\mathbf{PA}_{Y}$, $X\not\perp Y\mid\mathbf{PA}_{Y}\backslash\{X\}$ [57, p.109]. Accordingly, for $\mathcal{M}_{1}$, we show that both $(1)Y_{t+1}\not\perp U_{t}|A_{t}$ and $(2)Y_{t+1}\not\perp A_{t}|U_{t}$. All probabilities are calculated by summing over the relevant rows of Table 5.
  1. 1.
    $Y_{t+1}\not\perp U_{t}|A_{t}$
    $\mathbb{P}(Y_{t+1}=1,U_{t}=0|A_{t}=1)$ $=\frac{\mathbb{P}(Y_{t+1}=1,U_{t}=0,A_{t}=1)}{\mathbb{P}(A_{t}=1)}$
    $=\frac{p^{2}-p^{3}}{2p-p^{2}}$
    $=\frac{p^{2}-p}{p-2}$
    $\mathbb{P}(Y_{t+1}=1|A_{t}=1)\times\mathbb{P}(U_{t}=0|A_{t}=1)$ $=\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=1)}{\mathbb{P}(A_{t}=1)}\times\frac{\mathbb{P}(U_{t}=0,A_{t}=1)}{\mathbb{P}(A_{t}=1)}$
    $=\frac{-p^{3}+p^{2}+p}{2p-p^{2}}\times\frac{p-p^{2}}{2p-p^{2}}$
    $=\frac{p^{3}-2p^{2}+1}{p^{2}-4p+4}$
    Unless $p=1$ (ruled out by assumption), $\frac{p^{2}-p}{p-2}\neq\frac{p^{3}-2p^{2}+1}{p^{2}-4p+4}$. Thus, $\mathbb{P}(Y_{t+1}=1,U_{t}=0|A_{t}=1)\neq\mathbb{P}(Y_{t+1}=1|A_{t}=1)\times\mathbb{P}(U_{t}=0|A_{t}=1)$ and so $Y_{t+1}\not\perp U_{t}|A_{t}$.
  2. 2.
    $Y_{t+1}\not\perp A_{t}|U_{t}$
    $\mathbb{P}(Y_{t+1}=1,A_{t}=1|U_{t}=0)$ $=\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=1,U_{t}=0)}{\mathbb{P}(U_{t}=0)}$
    $=\frac{p^{2}-p^{3}}{1-p}$
    $=p^{2}$
    $\mathbb{P}(Y_{t+1}=1|U_{t}=0)\times\mathbb{P}(A_{t}=1|U_{t}=0)$ $=\frac{p^{2}-p^{3}}{1-p}\times\frac{p-p^{2}}{1-p}$
    $=p^{2}\times p$
    $=p^{3}$
    Unless $p=0$ or $p=1$ (ruled out by assumption), $p^{2}\neq p^{3}$. Thus $\mathbb{P}(Y_{t+1}=1,A_{t}=1|U_{t}=0)\neq\mathbb{P}(Y_{t+1}=1|U_{t}=0)\times\mathbb{P}(A_{t}=1|U_{t}=0)$ and so $Y_{t+1}\not\perp A_{t}|U_{t}$.
Meanwhile, $\mathcal{M}_{2}$ meets causal minimality vacuously. Since there are only two model variables, each variable has (at most) one parent.
Finally, we show $\mathcal{M}_{1}$ meets the three criteria of expertly targeted action (Definition 4.1) while $\mathcal{M}_{2}$ does not.
We begin by showing $\mathcal{M}_{1}$ meets all three criteria:
  1. 1.
    Effective action:
    $\mathbb{P}(Y_{t+1}|do(A_{t}=1),U_{t}=1)$ $=\mathbb{P}(Y_{t+1}|A_{t}=1,U_{t}=1)$ Parents do/see
    $=\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=1,U_{t}=1)}{\mathbb{P}(U_{t}=1,A_{t}=1)}$ Definition of conditional prob.
    $=\frac{p}{p}$ Table 5 and simplification
    $=1$
    $>0$
    $=\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=0,U_{t}=1)}{\mathbb{P}(A_{t}=0,U_{t}=1)}$ Table 5 and simplification
    $=\mathbb{P}(Y_{t+1}|A_{t}=0,U_{t}=1)$ Definition of conditional prob.
    $=\mathbb{P}(Y_{t+1}|do(A_{t}=0),U_{t}=1)$ Parents do/see
    Parents do/see is that conditioning and intervening on parent values of a variable are interchangeable [11, p.32]. It is true for all DAGs generated by SCMs [11, p.33].
  2. 2.
    Targeted action:
    $\mathbb{P}(A_{t}=1|do(U_{t}=1))$ $=\mathbb{P}(A_{t}=1|U_{t}=1)$ Parents do/see
    $=\frac{\mathbb{P}(A_{t}=1,U_{t}=1}{\mathbb{P}(U_{t}=1)}$
    $=\frac{p}{p}$ Table 5 and simplification
    $=1$
    $>p$ $p\in(0,1)\text{ by assumption }$
    $=\frac{p-p^{2}}{1-p}$ Algebraic manipulation
    $=\frac{\mathbb{P}(A_{t}=1,U_{t}=0)}{\mathbb{P}(U_{t}=0)}$ Table 5 and simplification
    $=\mathbb{P}(A_{t}=1|do(U_{t}=0))$ Definition of conditional prob. and parents do/see
  3. 3.
    Heterogeneous action:
    $\mathbb{P}(Y_{t}=1|do(A_{t}=1,U_{t}=1))$
    $-\mathbb{P}(Y_{t}=1|do(A_{t}=0,U_{t}=1))$ $=\mathbb{P}(Y_{t}=1|A_{t}=1,U_{t}=1)$
    $-\mathbb{P}(Y_{t}=1|A_{t}=0,U_{t}=1)$ Parents do/see
    $=\frac{\mathbb{P}(Y_{t}=1,A_{t}=1,U_{t}=1)}{\mathbb{P}(A_{t}=1,U_{t}=1)}$ Definition of conditional prob.
    $-\frac{\mathbb{P}(Y_{t}=1,A_{t}=0,U_{t}=1)}{\mathbb{P}(A_{t}=1,U_{t}=1)}$ Definition of conditional prob.
    $=\frac{p}{p}-\frac{0}{p}$
    $=1$
    $>p$ $p\in(0,1)$
    $=\frac{p^{2}-p^{3}}{p-p^{2}}-0$ Algebraic manipulation
    $=\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=1,U_{t}=0)}{\mathbb{P}(A_{t}=1,U_{t}=0)}$
    $-\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=0,U_{t}=0)}{\mathbb{P}(A_{t}=0,U_{t}=0)}$ Table 5 and simplification
    $=\mathbb{P}(Y_{t}=1|do(A_{t}=1,U_{t}=0))$
    $-\mathbb{P}(Y_{t}=1|do(A_{t}=0,U_{t}=0))$ Definition of conditional prob. and parents do/see
Meanwhile, note that $\mathcal{M}_{2}$ cannot meet targeted action nor heterogeneous action, since $U_{t}$ is not included in the model.
Thus, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are causally minimal models inducing equivalent observational distributions $\mathbb{P}(A_{t},Y_{t+1})$, yet only $\mathcal{M}_{1}$ exhibits expertly targeted action. ∎
A.2.2 Proof of Proposition 4.3
Proof.
By the Markov condition, if $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$ then $A_{t}\not\perp_{\mathcal{G}_{\mathcal{M}}}Y_{t+1}|\ \mathbf{X}_{t}$. Thus, by the definition of d-separation, either $\mathbf{U}_{t}\to A_{t}$ or $A_{t}\to Y_{t+1}$.
First, consider the case that $A_{t}\rightarrow Y_{t+1}$. Then, for some $X_{t}=x$, $A_{t}=a$, $U_{t}=u$ and $Y_{t+1}=y$ we have:
$\mathbb{P}(Y_{t+1}=1,A_{t}=a|X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=1|X_{t}=x,U_{t}=u)$
×$\quad\times\mathbb{P}(A_{t}=a|X_{t}=x,U_{t}=u)$ Causal minimality
$\mathbb{P}(Y_{t+1}=1|A_{t}=a,X_{t}=x,U_{t}=u)$
×$\quad\times\mathbb{P}(A_{t}=a|X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=1|X_{t}=x,U_{t}=u)$
×$\quad\times\mathbb{P}(A_{t}=a|X_{t}=x,U_{t}=u)$ Product rule
$\mathbb{P}(Y_{t+1}=1|A_{t}=a,X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=1|X_{t}=x,U_{t}=u)$ Cancellation
$\mathbb{P}(Y_{t+1}=y|A_{t}=a,X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(A_{t}=1)\times\mathbb{P}(Y_{t+1}=y|A_{t}=1,X_{t}=x,U_{t}=u)$
$\quad+\mathbb{P}(A_{t}=0)\times\mathbb{P}(Y_{t+1}=y|A_{t}=0,X_{t}=x,U_{t}=u)$ Law of total probability
$\mathbb{P}(Y_{t+1}=y|A_{t}=a,X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=y|A_{t}=0,X_{t}=x,U_{t}=u)$∗Simplification
$\mathbb{P}(Y_{t+1}=y|do(A_{t}=1),X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=y|do(A_{t}=0),X_{t}=x,U_{t}=u)$∗∗Parents do/see
$\mathbb{P}(Y_{t+1}=1|do(A_{t}=1),X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=1|do(A_{t}=0),X_{t}=x,U_{t}=u)$ $Y_{t+1}$ binary
Recall that causal minimality is equivalent to that, for all causal parents of $Y$, $X\in\mathbf{PA}_{Y}$, $X\not\perp Y\mid\mathbf{PA}_{Y}\backslash\{X\}$ [57, p.109].
∗∗ Suppose $a=1$. Then by subtraction from both sides,
$\mathbb{P}(A_{t}=0)\times\mathbb{P}(Y_{t+1}=y|A_{t}=1,X_{t}=x,U_{t}=u)\neq\mathbb{P}(A_{t}=0)\times\mathbb{P}(Y_{t+1}=y|A_{t}=0,X_{t}=x,U_{t}=u)$
And so by cancellation,
$\mathbb{P}(Y_{t+1}=y|A_{t}=1,X_{t}=x,U_{t}=u)\neq\mathbb{P}(Y_{t+1}=y|A_{t}=0,X_{t}=x,U_{t}=u)$
The reasoning is symmetrical if $a=0$.
∗∗∗ Recall parents do/see is that conditioning and intervening on parent values of a variable are interchangeable [11, p.32]. It is true for all DAGs generated by SCMs [11, p.33].
Thus $A_{t}\rightarrow Y_{t+1}$ implies impactful action is met. $U_{t}\rightarrow A_{t}$ implies non-algorithmic action is met by a very similar argument. Thus, since either $A_{t}\rightarrow Y_{t+1}$ or $U_{t}\rightarrow A_{t}$ or both, also either impactful action is met, non-algorithmic action is met, or both.
[…]
(1) Impactful action
[…]
(2) Non-algorithmic action
Figure 4: The two conditions of impactful action and non-algorithmic action are implied by certain causal edges being present in the graph. The proof of Proposition 4.3 can thus use d-separation and Markovity to show these criteria are met.
A.2.3 Generalization to Nonbinary Contrasts
Stating three causal conditions of expertly targeted action for nonbinary contrasts requires determining the importance of each causal contrast for each action and algorithmic unobservable valuation. For instance, in determining whether a particular action taken by an advisor (say, recommending a change to a student’s schedule in a meeting), was effective at improving a student’s outcome (say, increasing their chances of graduation), what other action should we contrast it to? Should we contrast it to taking some other action (say, recommending a student speak with their instructor)? If so, which one? Should we contrast it to taking no other action, imagining that the advisor had not met with the student at all? Or should we combine these (and other) possibilities together somehow?
This is ultimately a normative choice with multiple defensible answers, depending on what alternate possibilities we deem relevant to attributing expertise. One way to formalize potential choices would be to stipulate weights on the relative importance of each causal contrast. For instance, to generalize effective action, for each $a\in R_{A_{t}}$, the support of $A_{t}$, define a collection of importance weights on the constrasts $W_{a}=\{w_{b}:b\in R_{A_{t}}/\ a\}$ such that $\sum W_{a}=1$. A similar weighting can be defined for $\textbf{U}_{t}$, call it $W_{u}$. Then the causal contrasts to the action and non-algorithmic context evaluated in the components of the definition are weighted according to $W_{a}$ and $W_{u}$. For instance, effective action becomes checking that the particular action taken by the advisor, $a$, is more likely to have produced a good outcome than the alternative actions would have been, with the alternative actions weighted by $W_{a}$.
A.3 Empirical Analysis of Advisor Expertise
A.3.1 Test Population
Figure 6 visualizes the number of students enrolled each semester and Figure 6 visualizes the number of meetings per semester.
Figure 5: Enrolled students per semester
Figure 6: Number of meetings per semester
A.3.2 Positive vs. Negative Interventions
The ExpertTest* only rejects the null whenever the mean squared error in the synthetic datasets is routinely higher than the the mean squared error of the original dataset. In Alur et al. [5], this one-sided test of predictions is sensible, as it ensures the test only detects instances where non-algorithmic information is used to improve the predictions of the human forecaster.
However, in the setting of intervention, we wish to detect actions that negatively predict outcomes, too, as these might indicate targeting of interventions to at-risk students. Thus we might be curious about whether actions $A_{t}$ predict $1-Y_{i}$, formally captured by taking $MSE^{-}(\mathcal{D})=\sum_{i=1}^{n}(a_{t}^{i}-|1-y_{t+1}|)^{2}=\sum_{i=1}^{n}(|1-a_{t}^{i}|-y_{t+1})^{2}$ (for binary $A_{t}^{i},Y_{t+1}$). This corresponds to running two one-tailed tests on both directions of association.
Another approach would be to run a two-tailed test. Without loss of generality, suppose $\tau_{K}=\frac{J}{K}\leq 0.5$; then we would take the $p$-value as $\mathbb{P}(\tau_{K}\geq\frac{K-J}{K}\text{ or }\tau_{K}\leq\frac{J}{K})=2\times\frac{J+1}{K+1}$. While this test doubles uncorrected p-values, it halves the number of tested hypotheses; thus, post-correction, the results are quite similar. For instance, tests of persistence yield the same results of just Intervention 17 (Recommend Schedule) as significant. For the GPA outcome, all the same interventions are significant except Intervention 2 (Correct Schedule), which drops just below significance ($p=0.059$).
We opt for the one-tailed test with transformed actions because it is simpler to interpret. The test of $A_{t}^{+}$ shows whether actions predict positive outcomes. The test of $A_{t}^{-}$ shows whether inaction predicts positive outcomes, equivalent to actions predicting negative outcomes. We note that the two-tailed test can be interpreted in a similar way, but it requires algebraically manipulating the test statistic.
A.3.3 Test Versions and Parameters
The ExpertTest* requires two parameters: $L$, the number of pairs, and $K$, the number of iterations.
As mentioned in §5.1.2, we greedily pair $L$ “approximately-identical” students at baseline using a Euclidian distance function on fixed-scaled administrative features (excluding their race and ethnicity, which Georgia State University excludes from its predictors), prohibiting cross-semester matches for tests on semesterly outcomes. We select the number of pairs $L=100$ for cumulative tests, $L=800$ for semesterly GPA tests, and $L=700$ for semesterly persistence tests, to yield an acceptable tradeoff of the number of pairs (which increases test power) with the quality of the matches (the more imperfect the matches, the higher the test’s error).
For cumulative ExpertTest*, we visualize the differences by attribute between paired students in Figure 7 in the original units (e.g. date of birth in days, test scores in points). The maximum Euclidian distance between paired students in scaled feature space is $\approx 0.25$; when students are further away in one attribute, it is compensated for by a close match in other attributes. Pairs for the semesterly version of the test are slightly different (as there is a slightly different mix and smaller number of students enrolled each semester); Figure 7 is representative of match quality for these tests, with the highest Euclidian distance per semester between pairs rising at most to 0.49 in scaled feature space.
Figure 7: Differences between paired students’ attributes (excluding race and ethnicity) at baseline for $L=100$ pairs.
We set $K=1000$ to balance the trade-off between minimizing the ExpertTest* type I error—which decreases as $K$ increases [5]—with computational constraints. The test’s type I error includes an additive $\frac{1}{K+1}$ term, so the choice of $K=1000$ ensures that this component contributes no more than $0.001$ to the test’s error.
Additionally, we run the test on two timescales (early intervention and semesterly), and multiple different populations and outcome variables to ensure results are robust.
Table 7 summarizes all our test parameters, variables and exclusions.
Population $A_{t}$ $\mathbf{Y}_{t+1}$ $L$ $K$
Semesterly enrolled non-transfers* Term intervention $i$ for $i\in[1,20]$ Next-term persistence 700 1000
Semesterly enrolled non-transfers Term intervention $i$ for $i\in[1,20]$ Term GPA $\geq 2.75$ 800 1000
Semesterly enrolled non-transfers Term intervention $i$ for $i\in[1,20]$ Term GPA $\geq 3.0$ 800 1000
Semesterly enrolled non-transfers Term intervention $i$ for $i\in[1,20]$ Term GPA $\geq 3.25$ 800 1000
Semesterly enrolled non-transfers Term intervention $i$ for $i\in[1,20]$ Term GPA $\geq 3.5$ 800 1000
Semesterly enrolled non-transfers Term intervention $i$ for $i\in[1,20]$ Term GPA $\geq 3.75$ 800 1000
Semesterly enrolled non-transfers Term intervention $i$ for $i\in[1,20]$ Term GPA $\geq 4.0$ 800 1000
Semesterly enrolled all transfer status* Term intervention $i$ for $i\in[1,20]$ Next-term persistence 700 1000
Semesterly enrolled all transfer statuses Term intervention $i$ for $i\in[1,20]$ GPA $\geq 3.5$ 800 1000
Semesterly enrolled non-transfer non-advisor 1* Term intervention $i$ for $i\in[1,20]$ Next-term persistence 700 1000
Semesterly enrolled non-transfer non-advisor 1 Term intervention $i$ for $i\in[1,20]$ GPA $\geq 3.5$ 800 1000
Enrolled first year non-transfer First-year intervention $i$ for $i\in[1,20]$ 4-year graduation 100 1000
Enrolled first year non-transfer First-year intervention $i$ for $i\in[1,20]$ Cumulative GPA $\geq 3.5$ 100 1000
Table 7: Versions of the ExpertTest* we run with different populations, $A_{t}$, $\mathbf{Y}_{t+1}$, $L$ (the number of pairs), and $K$ (the number of synthetic datasets).
*For tests on next-term persistence, we exclude enrolled students from spring 2020 in the population as the experiment concluded that term and no subsequent persistence data is available.
A.3.4 Number of Swaps and Test Power
Table 8 shows the number of swaps between algorithmically-indistinguishable students that could increase or decrease the mean squared error (MSE) of synthetic datasets, for the semesterly tests on both persistence and GPA $\geq 3.5$ outcomes excluding potential transfer students. A smaller number of such swaps—whether increasing or decreasing the MSE—reduces the power of the test to detect conditional dependence. This is because when few swaps change the MSE, synthetic datasets are less likely to differ from the true dataset in MSE, even if non-algorithmic information is being used.
Thus, test power tends to decrease as the base rate of either the intervention or the outcome approaches 0 or 1. When interventions are rarely applied, most students will have the same intervention status, leaving little room for swaps to change the MSE. Similarly, because the base rate of persistence is relatively high (see Figure 6), students are less likely to differ on this outcome, which reduces the test’s power for persistence compared to GPA outcomes.
Intervention Count $\downarrow$ MSE GPA $\uparrow$ MSE GPA $\downarrow$ MSE Persistence $\uparrow$ MSE Persistence
1. Seek academic support 162 19 11 11 5
2. Correct schedule 503 55 32 14 11
3. Declare or change major 238 23 18 8 4
4. Progression issues 666 63 35 20 12
5. Adjust course balance 192 22 19 11 10
6. See instructor 73 10 6 3 0
7. Seek academic coaching 14 1 0 0 0
8. Seek departmental or faculty guidance 190 15 23 6 3
9. Address financial issues 338 25 31 9 10
10. Address non-financial collegiate requirements 280 23 26 6 8
11. Seek career counseling and service 307 23 31 5 11
12. Seek engagement opportunities on campus 253 13 29 6 13
13. See professional counselor 28 1 4 2 2
14. Seek disability services on campus 2 0 0 0 0
15. See medical professional 9 1 0 0 0
16. Schedule an appointment 1035 76 45 25 13
17. Recommended schedule 1601 50 66 8 24
18. See primary program advisor 49 6 7 1 3
19. Academic map/degree planner 635 7 19 14 6
20. Other (Describe in comments) 1810 41 70 10 23
Table 8: At $L=800$ pairs for GPA outcomes and $L=700$ for persistence outcomes, the above table records the total number of swaps of intervention $i$ for approximately algorithmically-indistinguishable pairs that could decrease ()($\downarrow$)( ↓ ) vs. increase ()($\uparrow$)( ↑ ) the mean squared error of the intervention at predicting the outcomes. Counts of the number of times the intervention was performed total are also included.
A.3.5 Full Results
Results of the ExpertTest* are reported in Table 9 for term tests and Table 10 for early-intervention tests. Additionally, for term tests (which were the only tests with significant results), results including potential transfer students are reported in Table 11 and excluding students assigned Advisor 1 in Table 12. Results are unchanged by exclusion of advisor 1, and remain similar upon inclusion of potential transfer students, which adds a new intervention, Intervention 8 (Seek Department or Faculty Guidance) as positively predictive of a high term GPA, and removes Intervention 2 (Correct Schedule) as negatively predictive of a low term GPA. The slight differences are likely reflective of slightly different characteristics of transfer student population and the inherent randomness of the test (small differences in the generated synthetic datasets).
Additionally, the top 5 most significant interventions at GPA binarization thresholds ranging from 2.5 to 4.0 are included in Table 13. As the table demonstrates, Interventions 4 (Progression Issues) and 16 (Schedule an Appointment) are significant for all tests, while Intervention 2 (Correct Schedule) and Intervention 20 (Other) are always in the top 5 except for in predicting a perfect GPA of 4.0 (which would be expected due to the especially high bar in addition to the reduced power of the test when most people’s outcomes do not differ).
Intervention Term GPA (+)(+)( + ) Term GPA (+)(+)( + ) Next-term persistence (+)(+)( + ) Next-term persistence ()(-)( - )
1 0.998002 0.399600 1.000000 0.399600
2 0.998002 0.049950* 1.000000 0.882451
3 0.998002 0.633652 1.000000 0.607393
4 0.998002 0.049950* 1.000000 0.607393
5 0.998002 0.777001 0.923692 0.882451
6 0.998002 0.519481 1.000000 0.399600
7 0.998002 0.639361 0.882451 0.882451
8 0.444000 0.998002 1.000000 0.624830
9 0.633652 0.998002 0.882451 0.923692
10 0.777001 0.998002 0.882451 1.000000
11 0.515848 0.998002 0.399600 1.000000
12 0.053280 0.998002 0.399600 1.000000
13 0.455544 0.998002 0.882451 0.882451
14 0.970458 0.961039 0.882451 0.882451
15 0.998002 0.639361 0.882451 0.882451
16 0.998002 0.039960* 1.000000 0.333000
17 0.399600 0.998002 0.039960* 1.000000
18 0.832851 0.998002 0.715951 1.000000
19 0.053280 0.998002 1.000000 0.399600
20 0.039960* 0.998002 0.119880 1.000000
Table 9: Results of ExpertTest* on term outcomes excluding data from students missing data on transfer status, with Benjamini-Hochberg correction applied.
Intervention Persistence (+)(+)( + ) Persistence ()(-)( - ) Graduation (+)(+)( + ) Graduation ()(-)( - )
1 0.866825 0.948143 0.943672 0.839161
2 0.864590 0.948143 0.943672 0.839161
3 0.864590 0.955801 0.839161 0.943672
4 0.948143 0.880786 0.936563 0.839161
5 0.866825 0.948143 0.943672 0.839161
6 0.949051 0.864590 0.875791 0.875791
7 0.880786 0.880786 0.875791 0.875791
8 0.948143 0.880786 0.839161 0.943672
9 0.864590 0.948143 0.839161 0.936563
10 0.880786 0.948143 0.839161 0.943672
11 0.346320 0.993007 0.839161 0.936563
12 0.864590 0.949051 0.839161 0.936563
13 0.880786 0.880786 0.936563 0.839161
14 0.880786 0.880786 0.875791 0.875791
15 0.948143 0.864590 0.875791 0.875791
16 0.948143 0.864590 0.966034 0.839161
17 0.239760 0.993007 0.839161 0.943672
18 0.239760 0.993007 0.875791 0.936563
19 0.880786 0.880786 0.936563 0.875791
20 0.864590 0.949051 0.839161 0.936563
Table 10: Results of ExpertTest* on early interventions for 4-year outcomes, excluding data from students missing data on transfer status, with Benjamini-Hochberg correction applied.
Intervention Next-term persistence (+)(+)( + ) Next-term persistence ()(-)( - ) Term GPA (+)(+)( + ) Term GPA ()(-)( - )
1 1.000000 0.446220 1.000000 0.049950
2 0.931925 0.882451 1.000000 0.164835
3 0.882451 0.931925 1.000000 0.575425
4 0.931925 0.882451 1.000000 0.204240
5 0.931925 0.882451 1.000000 0.794994
6 1.000000 0.693852 0.817183 1.000000
7 0.882451 0.882451 1.000000 0.579421
8 1.000000 0.446220 0.049950 1.000000
9 0.446220 1.000000 0.399600 1.000000
10 0.722611 1.000000 0.485668 1.000000
11 0.446220 1.000000 0.399600 1.000000
12 0.446220 1.000000 0.063936 1.000000
13 0.882451 0.882451 0.399600 1.000000
14 0.882451 0.882451 0.579421 1.000000
15 0.882451 0.882451 1.000000 0.579421
16 1.000000 0.259740 1.000000 0.039960
17 0.019980 1.000000 0.108463 1.000000
18 0.693852 1.000000 0.575425 1.000000
19 1.000000 0.066600 0.093240 1.000000
20 0.019980 1.000000 0.039960 1.000000
Table 11: Results of ExpertTest* on semesterly interventions, including students with missing transfer status, with Benjamini-Hochberg correction applied.
Intervention Next-term persistence (+)(+)( + ) Next-term persistence ()(-)( - ) Term GPA (+)(+)( + ) Term GPA ()(-)( - )
1 0.998002 0.365349 0.998002 0.399600
2 0.994903 0.775695 0.998002 0.049950
3 0.994903 0.775695 0.998002 0.633652
4 0.994903 0.365349 0.998002 0.049950
5 0.994903 0.775695 0.998002 0.777001
6 0.994903 0.775695 0.998002 0.519481
7 0.920818 0.920818 0.998002 0.639361
8 0.994903 0.365349 0.444000 0.998002
9 0.994903 0.775695 0.633652 0.998002
10 0.775695 0.994903 0.777001 0.998002
11 0.775695 0.994903 0.515848 0.998002
12 0.365349 0.994903 0.053280 0.998002
13 0.994903 0.775695 0.455544 0.998002
14 0.920818 0.920818 0.970458 0.961039
15 0.920818 0.920818 0.998002 0.639361
16 0.994903 0.365349 0.998002 0.039960
17 0.159840 0.998002 0.399600 0.998002
18 0.775695 0.994903 0.832851 0.998002
19 0.994903 0.775695 0.053280 0.998002
20 0.365349 0.994903 0.039960 0.998002
Table 12: Results of ExpertTest* on semesterly interventions, excluding data from students missing data on transfer status, and excluding Advisor 1, with Benjamini-Hochberg correction applied.
Term GPA $\mathbf{\geq 2.75}$ $\mathbf{\geq 3.0}$ $\mathbf{\geq 3.25}$ $\mathbf{\geq 3.75}$ $\mathbf{\geq 4.0}$
4 (-) $p=0.019980^{*}$ 4 (-) $p=0.019980^{*}$ 4 (-) $p=0.019980^{*}$ 4 (-) $p=0.019980^{*}$ 4 (-) $p=0.019980^{*}$
16 (-) $p=0.019980^{*}$ 16 (-) $p=0.019980^{*}$ 16 (-) $p=0.019980^{*}$ 16 (-) $p=0.019980^{*}$ 16 (-) $p=0.019980^{*}$
2 (-) $p=0.026640^{*}$ 12 (+) $p=0.053280$ 2 (-) $p=0.026640^{*}$ 17 (+) $p=0.303696$ 6 (-) $p=0.346320$
12 (+) $p=0.039960^{*}$ 20 (+) $p=0.055944$ 12 (+) $p=0.069930$ 20 (+) $p=0.303696$ 4 (-) $p=0.499500$
20 (+) $p=0.446220$ 2 (-) $p=0.055944$ 9 (-) $0.139860$ 2 (-) $p=0.303696$ 19 (+) $0.511489$
Table 13: Top five most significant interventions on the ExpertTest* at GPA thresholds ranging from 2.75 to 4.0, and their $p$-values with Benjamini-Hochberg correction applied.
A.3.6 Correlations Between Interventions
A significant result on the ExpertTest* might reflect dependence driven by application of interventions together, since advisors may apply multiple interventions in one meeting. To address this possibility without having to test $2^{20}$ hypotheses (every possible combination of interventions), we here interpret results in light of the Pearson correlations between significant interventions, visualized in Figure 8, on the assumption that any interventions that may be significant due to joint application will be detected as separately significant.
For persistence outcomes, only one intervention (Intervention 17, Recommend Schedule) is significant, so no correlations warrant further attention. For GPA outcomes, Intervention 4 (progression issues) shows moderate correlation with each of the other significant interventions: Intervention 2 (Correct Schedule, $r=0.44$, $p=6.4\times 10^{-117}$), Intervention 16 (Schedule Appointment, $r=0.35$, $p=1.31\times 10^{-72}$), and Intervention 20 (Other, $r=0.30$, $p=1.5\times 10^{-53}$), while the remaining pairs of interventions show only weak correlation. Notably, Intervention 20 (Other) predicts a GPA $\geq 3.5$ positively while Intervention 4 (Progression Issues) predicts it negatively, making it unlikely that their conditional dependence with outcomes reflects joint causal effects or targeting. Meanwhile, qualitative analysis supports the interpretation that Interventions 2 (Correct Schedule) and 16 (Schedule Appointment) are informed by different types of non-algorithmic information than Intervention 4 (Progression Issues) (Table 3), suggesting they are separately targeted. However, given the moderate correlations, it remains possible that Interventions 2 (Correct Schedule) and 4 (Progression Issues), as well as Interventions 4 (Progression Issues) and 16 (Schedule Appointment), are either jointly non-algorithmically targeted or jointly influence GPA outcomes.
Figure 8: Pearson correlation matrix between intervention per semester variables $A_{t}^{i}$ for $i\in[1,20]$.
A.4 Exploratory Qualitative Analysis
A.4.1 Codebook for the Exploratory Qualitative Analysis.
In our analysis, codes served as succinct, essence-capturing labels that represent meaningful patterns in the data [67]. We organized codes hierarchically, with high-level categories capturing key differences in advisor style and aims (e.g. discussing a topic open-endedly, making directive suggestions to the student, or building rapport with the student) and finer-grained codes differentiating interactions within these categories (for instance, the discussion topic). We report the frequency of these applied codes by each advisor in Table 14. Our complete codebook with representative quotes is provided below.
  • Academic Discussion: Do the advisor and student discuss an academic topic (without advisor making a recommendation)?
    • Scheduling - discussing course schedules, registration timelines, or planning for upcoming semesters:
      “Student wants to change their schedule for the spring and take some courses online. We discussed several available options."
    • Academic administrative - addressing administrative issues like degree requirements, deadlines, navigating transfers, or institutional processes:
      “Reminded a student about the deadline to submit AP, DE, and/or IB transcripts."
    • Major/minor selection - discussing the student choice of major/minor:
      “Student is thinking about changing majors to finance. They like numbers, but also enjoy being creative. I gave them the Focus 2 Assessment to help identify their strengths."
    • Academic performance - reviewing grades, progress, or challenges in coursework:
      “Student wanted to withdraw from [ECON XXX], as their first exam was a 40%; they admitted they just didn’t study. We talked through their degree requirements and they understand the importance of taking this class seriously now."
    • EAB/algorithmic review - reviewing algorithmically-suggested or institutional benchmarks for the student, like EAB’s Success Markers or Academic Improvement Plan warnings:
      “I have printed and reviewed degree audit, reviewed EAB for success markers, and notified student.”
  • Nonacademic Discussion: Do the advisor and student discuss a non-academic topic (without advisor making a recommendation)?
    • Career - talking about career aspirations, opportunities, career fairs, or preparation:
      “Student is interested in going to law school. I explained popular majors for pre-law route such as POLS, IDS, and Philosophy.”
    • Financial aid - discussing applying for or maintaining financial aid packages/status:
      “Student mentioned they will go to financial aid due to their refund not going to the right bank. This is vital as they have not been able to get their school books yet."
    • Financial trouble - discussing problems students are having with finances:
      “Student is on the drop-list for non payment. I called the student to warn them and to come up with a plan, and talk about reinstatement."
    • Other nonacademic - discussing any other nonacademic matter, such as extracurriculars, housing/commuting, or family, medical concerns:
      “Student is living on campus, has joined several social organizations, and we talked about work-life balance."
  • Academic Suggestion: How is the advisor guiding the student academically?
    • Follow up with advisor - the advisor asks students to check in with them, schedule a meeting with them, or follow up in the future with updates about a specifically academic matter:
      “Asked student to sit down together and discuss each hold, and help them figure out where to go to get them resolved."
    • Follow up with another person - the advisor asks students to speak with another person about a specifically academic matter:
      “I encouraged them to talk to their instructors of [philosophy course] and [film course], to ask about any tips for the remaining of the semester."
    • Guiding/steering the student - the advisor tells a student how to alter their program of study, for instance by changing major, recommending a schedule, or adding/dropping a class:
      “I told them I’m nervous about them having GEOG and MATH together and suggested to revise the schedule; however, they were persistent about completing them both in the Spring semester."
  • Nonacademic Suggestion: How is the advisor guiding the student non-academically?
    • Follow up with advisor - the advisor asks students to check in with them later on or lets them know that they can check in with future questions about specifically nonacademic matters:
      “Student will return in the near future with top 5 schools they are considering for medical school."
    • Other suggestion - the advisor makes any other sort of suggestion about specifically non-academic matters, for instance referring students to a campus office or recommending a student pursue an extracurricular:
      “Advisor confirmed that student can still fulfill law school requirements with courses and completed what-if analysis. Advised student to visit career services for further confirmation and to complete assessments. "
  • Generic Suggestion: How is the advisor guiding the student across both academic/nonacademic matters?
    • Follow up with advisor - the advisor asks students to check in with them later / follow up later, without mentioning any specific topic:
      “Student will continue to explore their options as far as choosing a major. Asked the student to follow up if they have any additional questions or concerns."
  • Relationship: How does the advisor get to know/support the student emotionally?
    • Empathy/sympathy - expressing concern, understanding, or emotional support for the student’s circumstances or challenges:
      “I am very proud of you for completing your 1st year of college and looking forward to your success in the upcoming academic year."
    • Student affect - observing and noting the student’s emotional state or behavior during the interaction:
      “Student really likes GSU, joined Greek life. I asked about their school and social life balance was going, they said good."
  • Intervention: Does the advisor take action on behalf of the student?
    • Changing the student’s schedule - directly altering a student’s schedule on their behalf:
      “I looked through student’s degree planner and their schedule, and adjusted it due to needing [MATH XXX]."
    • Contacting student’s parents - communicating directly with a student’s family members to address concerns or provide important information:
      “Spoke with student’s parent explaining to them that their kid needs to complete their FAFSA for the 17/18 academic year."
    • Speaking with department/instructor/administrator - The advisor reaches out to a student’s department, instructor, or an administrator on behalf of them.
      “Student had an issue with one of the professors. I walked them over to the Ombudsperson office to meet with Assistant Ombudsperson […] to discuss the issue."
    • Signing / filling out form for student - The advisor signs or completes a form on behalf of the student:
      “Student is within transient guidelines as acknowledged on the document. I completed the form, signed, and sent it to the registrar’s office for further processing."
Code Category Sub-code Advisor 1 Advisor 2 Advisor 3
Academic Discussion Scheduling 750 337 244
Academic administrative 298 90 28
Major/Minor Selection 214 97 80
Academic performance 190 33 38
EAB/algorithmic review 495 1 4
Academic Suggestion Follow up with advisor 185 13 19
Follow up with another person 52 29 23
Guiding/steering the student 115 51 129
Nonacademic Discussion Career 62 27 50
Financial aid 20 6 8
Financial trouble 33 19 24
Other nonacademic 26 7 32
Nonacademic Suggestion Follow up with advisor 39 0 7
Other suggestion 52 7 20
Generic Suggestion Follow up with advisor 642 1 60
Relationship Empathy/sympathy 42 3 20
Student affect 40 16 94
Intervention Changing student’s schedule 1 0 2
Contacting student’s parents 7 0 1
Speaking with department/instructor 2 2 1
Signing the form for student 24 15 2
Table 14: Frequencies of the applied codes by each advisor during student-advising meetings, based on our exploratory qualitative analysis. Notable differences in code distributions across advisors highlight distinct advising styles.
A.5 Qualitative Analysis of Advisor Expertise
A.5.1 Codebook for Qualitative Analysis of Non-Algorithmic Information
Below we include our codebook for types of reasons that advisors took interventions.
  • Course credits and placement: Matters to do with course placement and credit, including the student having prior credit from AP/dual enrollment, or discussion of the student’s performance on a placement exam for courses.
  • Course problem: The student is enrolled in the wrong course for their skill level, wants to drop/withdraw from a course they dislike, or is taking a course in which they are performing poorly.
  • Degree progress: The student is behind with degree progression.
  • Documents: The student has not turned in or checked their financial documents.
  • Financial regjstration problem: The student has a non-payment issue or is on the drop-list due to financial holds/reinstatements.
  • Load management: The student has too light or too heavy a course/extracurricular load.
  • Major choice: The student expresses preferences or indecision about major selection, often intertwined with discussion of career goals.
  • Non-financial registration problem: The student is unable to register due to having various non-financial holds on their account.
  • Personal circumstances: The student has various personal circumstances that are affecting their academic goals (health problems, family hardship, transportation considerations).
  • Low student engagement: The student did not attend their scheduled appointment, does not have an appointment, or is being standoffish in their current appointment.
  • Scholarship process: The student has a financial aid consideration, such as the student not meeting course requirements for their financial aid, or the student’s financial aid not having been posted to their account.
  • Student-initiated: The student explicitly asks the advisor for assistance or advice about an intervention.
  • Transfer/dropout: The student is considering transferring from or dropping out of Georgia State University.
  • Unregistered: The student is not registered for courses for the coming semester.
We identified the following types of “Other” interventions with $\geq 10$ instances of application:
  • Career planning: The advisor provides direct advice to the student on career or graduate school goals (rather than directing a student to career services, faculty, or student organizations). This includes administering career personality tests or helping students assemble a list of target graduate programs.
  • MAAPS information: The advisor explains the MAAPS experiment to students.
  • Placement test: The advisor provides counsel about placement tests to students, for instance advising a student to take a particular placement test.
  • Technical assistance: The advisor provides technical assistance to the student, such as teaching them how to register for courses.
  • Transfer and dual enrollment advising: The advisor provides advising to the student about enrolling at other institutions, including students considering transfers, dual enrollment, transient (semester) enrollment at another institution, or studying abroad.
We also identified the following types of “Other” interventions with $<10$ instances of application:
  • Academic improvement plan: The advisor explains the requirements of an academic improvement plan to students.
  • Administrative process assistance: The advisor helps the student with an administrative process, like petitioning for enrollment exceptions or filing administrative complaints.
  • Housing advice: The advisor provides advice on the student’s housing situation.
  • Online courses: The advisor suggests to the student that they should register for online courses.
  • Scholarship opportunities: The advisor points the student to relevant scholarship opportunities
We do not report results for these interventions since it is difficult to confirm that advisors intended to use the “Other” box to indicate application of these interventions without multiple examples.
A.5.2 Comments Mentioning Reasons
Table 15 records the number of meetings for all interventions (after exclusion of potential transfer students), as well as the number of comments associated with these meetings, the number of times the intervention was mentioned in these meetings, and the number of times the reason was mentioned for taking the intervention.
Intervention Sig. # Meetings # Comments # Intervention Mentioned # Reason Mentioned
2 (Correct Schedule) 756 64 38 36
4 (Progression Issues) 972 58 27 27
5 (Course Balance) $\times$ 241 35 15 14
9 (Fin. Issues) $\times$ 626 92 65 65
10 (Non-Fin. Requirements) $\times$ 348 13 3 3
16 (Schedule Appointment) 1343 347 254 157
17 (Recommend Schedule) 1915 791 486 243
20 (Other) 2384 484 189∗∗ 157∗∗
Table 15: Of the significant interventions and the three most insignificant interventions with a substantial number of meetings, counts of how many meetings with that intervention applied have comments left by advisors, comments that actually mention the relevant intervention, and comments that mention the reason(s) for taking that intervention. We can only analyze comments mentioning the reasons for taking an intervention for signs of non-algorithmic, outcome-relevant information.
See Appendix A.5.3 for details about customization vs. reasoning mentions.
∗∗ See Appendix A.5.4 for breakdown by intervention subtype.
A.5.3 Results for Intervention 17 (Recommend Schedule)
Without exception, when the advisors mention a reason for recommending a schedule, it is that the student had requested to discuss their schedule. In addition to the 243 mentions of a reason for schedule recommendations, there are 81 instances of advisors customizing recommended schedules to the student. Common customizations include the advisor recommending a specific class to a student based on the course placement test, the advisor suggesting schedule adjustments if the student has a heavy course load or is struggling in a particular class, and the advisor advising a student who is behind on degree requirements to modify their schedule to prioritize required courses and catch up. We note that customizations are likely a vast undercount, as we stringently required explicit evidence linking the information to their schedule recommendations that advisors do not often provide in their comments (for instance, simply mentioning the student wants to be a doctor and listing biology courses would not count; the advisor would have to explicitly say that the courses fulfilled pre-medical requirements).
A.5.4 Results for Intervention 20 (Other)
For Intervention 20 (Other), we identified five common types of interventions advisors took on comments labeled “Other,” listed with counts in Table 16. Without exception, the reasons for “Other” interventions (when they are mentioned) are pieces of information about the student that closely parallel the type of “Other” intervention itself; for instance, transfer/dual enrollment advising, such as helping students plan their schedule to facilitate a transfer, were all informed by learning about a student’s particular transfer or dual enrollment aspirations or situation.
“Other” Intervention # Interventions # Reasons
Career advising 45 42
MAAPS information 20 0
Placement test 57 53
Technical assistance 10 6
Transfer and dual credits 44 43
Table 16: For meetings with comments where advisors indicated they applied Intervention 20 (Other) we provide counts of mentions of the intervention and reasons in the comments, broken down by types of “Other” interventions with at least 10 instances of application.
A.5.5 Comparison to Insignificant Interventions
Unlike the significant interventions, which incorporate diverse, non-algorithmic information, two of the three insignificant interventions, Intervention 9 (Address Financial Issues) and Intervention 10 (Address Non-financial Collegiate Requirements), are dominated by administrative holds on registration, most commonly due to nonpayment of a balance on the student’s account. These administrative holds are plausibly either predictable from administrative data (for instance, financial holds might be predictable from students’ expected financial contribution) or not outcome-relevant (when the holds are resolved).
Intervention 5 (Adjust Course Balance) was an exception; though insignificant, advisors describe taking the intervention for similar reasons as Intervention 2 (Correct Schedule) and Intervention 4 (Progression Issues). We speculate that the insignificant result might be due to inconsistent use of this label given overlap in its meaning with these interventions. 75% of the 272 meetings in which advisors indicated applying Intervention 5 (Adjust Course Balance) also indicate applying Intervention 2 (Correct Schedule) or Intervention 4 (Progression Issues), and in our coding, the content of the comments appears quite similar (whether or not intervention 2 or 4 were also checked).
A.6 Quantitative investigation of heterogeneity in student outcomes
A.6.1 Exclusions
For all analyses, we standardize numerical variables and exclude students missing race and ethnicity data, transfer status, and cumulative GPA. Following the methodology in Rossman et al. [65], we convert composite SAT scores to ACT scores using a concordance table provided by the Educational Testing Service and ACT Inc. to facilitate uniform comparisons of test scores [1].
A.6.2 Heterogeneity by Assigned Advisor
We further note that only students in the treatment group are assigned to one of the MAAPS advisors; the students in the control group are assigned to a standard advisor. Thus, to look for heterogeneous treatment effects of the advisors, we fit the model (1) only to the treatment students, hence the lack of the main treatment effect term $T_{i}$. In addition to the above model, we also run a causal forest model [7] to capture potentially nonlinear heterogeneous treatment effects. We use 5-fold cross-validation and optimize the model to maximize the pure parameter heterogeneity score using the "het" criterion. Further methodological details and results are provided in Appendix A.6.4.
A.6.3 Heterogeneity by Student Demographics
Interaction terms between treatment $T_{i}$ and race-ethnicity in Equation 2 indicators are included to explore heterogeneous effects. In this analysis, students in the White group serve as the comparison class, meaning that the coefficients for the race-ethnicity indicators ($\beta_{2},\beta_{3},\beta_{4}$), and their interactions with treatment ($\beta_{5},\beta_{6},\beta_{7}$) represent deviations from the baseline effect observed for White students. We exclude with multi-racial race/ethnicity designations due to the small population size of this group. Similar to the analysis of heterogeneity by assigned advisor, we also run a causal forest to check for heterogeneity by student demographics.
A.6.4 Heterogeneity Analysis Using Causal Forests
We employ causal forests to model potential nonlinear heterogeneous treatment effects. This technique, based on generalized random forests [7], uses double machine learning (DML) to residualize treatment and outcome variables.
We specify the parameters below to ensure that the causal forest is flexible and uses the high out-of-sample predictive power. The tuned parameters directly impact the model’s ability to solve the local moment equation problem effectively:
$\begin{split}\mathbb{E}\left[\left(Y-\mathbb{E}[Y\mid X,W]-\langle\theta(x),T-\mathbb{E}[T\mid X,W]\rangle-\beta(x)\right)(T;1)\mid X=x\right]=0\end{split}$ (3)
where:
  • $X$ are the covariates, and $W$ are the control variables.
  • $\mathbb{E}[Y\mid X,W]$ and $\mathbb{E}[T\mid X,W]$ are fitted in the first stage using a cross-fitting procedure.
  • $\theta(x)$ represents the conditional average treatment effect at $X=x$ that we are trying to estimate.
  • The parameters $\beta(x)$ (intercept term) is iteratively refined during the tuning process.
We use 5-fold cross validation during the tuning process to select the optimal hyperparameters such as n_estimators (number of trees), max_depth (maximum depth of trees), and min_samples_split (minimum samples per split), and the native tune function from the CausalForestDML package for the causal forest:
  • The process tunes the major hyperparameters of the causal forest’s final stage by evaluating out-of-sample R-score performance: small forests of 100 trees are trained across a grid of hyperparameter values. Out-of-sample R-score performance is used to assess each configuration, after which all parameters are updated to their optimal values.
A.6.5 Power Analysis
To assess whether our dataset is sufficiently powered to detect meaningful HTEs, we employ a basic linear probability model. Our model estimates the effect of treatment on graduation outcomes while allowing for heterogeneity across particular racial groups. Specifically, we estimate the following equation:
$\mathbb{P}(Z_{i}=1)=\beta_{0}+\beta_{1}T_{i}+\beta_{2}\text{({Subgroup})}_{i}+\beta_{3}T_{i}\cdot\text{({Subgroup})}_{i}$ (4)
where Subgroup_i indicates the i-th student belonging to the chosen subgroup (e.g. race-ethnicity, gender, or assigned advisor).
Using a standard two-tailed hypothesis test at significance level $\alpha$ and power level $\beta$, we reject the null hypothesis if the test statistic $\left|\frac{\hat{\beta}_{3}}{{SE(\hat{\beta}_{3})}}\right|$ exceeds ($Z_{1-\alpha/2}$) in absolute terms. To achieve the desired power $Z_{1-\beta}$, we need $\frac{MDE}{{SE(\hat{\beta}_{3})}}\geq Z_{1-\alpha/2}+Z_{1-\beta}$, where $MDE$ is the smallest true effect of $\beta_{3}$. We estimate standard error of the empirical $\hat{\beta_{3}}$ distribution from our model and compute the minimum detectable effect (MDE) for this coefficient as follows:
$\begin{split}\text{MDE}=(Z_{1-\alpha/2}+Z_{1-\beta})\cdot\text{SE}(\hat{\beta}_{3})\end{split}$ (5)
where $Z_{1-\alpha/2}$ is the critical value from the standard normal distribution (1.96 for a 95% confidence level), $Z_{1-\beta}$ is the standard normal critical value for the desired power (0.84 for 80% power) and $\text{SE}(\hat{\beta}_{3})$ is the standard error of the estimated coefficient $\hat{\beta}_{3}$. We derive $\text{SE}(\hat{\beta}_{3})$ from the variance-covariance matrix of our model (4):
$\begin{split}&\text{SE}(\hat{\beta}_{3})=\sqrt{\text{Var}(\hat{\beta}_{3})}\\ \text{Var}(\hat{\beta}_{3})&=\sigma^{2}\left((X^{\prime}X)_{33}^{-1}\right)=\frac{\sum_{i=1}^{N}\hat{\epsilon}_{i}^{2}}{N-p}\left((X^{\prime}X)_{33}^{-1}\right)\end{split}$ (6)
where $N$ is the sample size, $\sigma^{2}$ is the variance of the error term $\hat{\epsilon}$ for the estimated parameters $p$, and $\left((X^{\prime}X)^{-1}\right)_{33}$ is the element corresponding to the coefficient $\hat{\beta}_{3}$ in the inverse of the Gram matrix $X^{\prime}X$ for the design matrix $X_{Nxp}$.
We further consider several experimental scenarios (e.g. different proportion of the subgroups and effect sizes) to analyze the minimum required sample size in order to achieve 80% power when detecting those effects. For these synthetic experiments, we calculate the coefficient $\hat{\beta}_{3}$ using the Differences-in-Differences (DiD) formula, which allows us to isolate the interaction effect between two subgroups:
$\begin{split}\hat{\beta}_{3}&=\Big{(}E[Z\mid T=1,S=1]-E[Z\mid T=0,S=1]\Big{)}-\Big{(}E[Z\mid T=1,S=0]-E[Z\mid T=0,S=0]\Big{)}\end{split}$ (7)
where $E[Z\mid T,S]$ represents the expected value of the outcome $Z$ given treatment status $T$ and subgroup membership $S$. Now, to achieve the objective of solving for the minimum required sample size N, we can rewrite equation (5) as follows:
$\begin{split}\text{MDE}^{2}&=(Z_{1-\alpha/2}+Z_{1-\beta})^{2}\cdot\left(\frac{\sum_{i=1}^{N}\hat{\epsilon}_{i}^{2}}{N-p}\right)\cdot(X^{\prime}X)_{33}^{-1}\\ &\frac{\sum_{i=1}^{N}\hat{\epsilon}_{i}^{2}}{N-p}=\frac{\text{MDE}^{2}}{(Z_{1-\alpha/2}+Z_{1-\beta})^{2}\cdot(X^{\prime}X)_{33}^{-1}}\\ &N=p+\frac{\sum_{i=1}^{N}\hat{\epsilon}_{i}^{2}\cdot(Z_{1-\alpha/2}+Z_{1-\beta})^{2}\cdot(X^{\prime}X)_{33}^{-1}}{\text{MDE}^{2}}\end{split}$ (8)
It is important to note that $SE(\hat{\beta}_{3})$ is inversely proportional to $\sqrt{N}$, which indicates that the sample size N is inversely proportional to the MDE, meaning that as the desired MDE decreases, the required sample size increases:
$\text{MDE}\propto\frac{1}{\sqrt{N}}$ (9)
Guided by the relation above, we show the computed MDE and the minimum required sample size N for various scenarios described by the model (4) next.
A.6.6 Power Limitations in Detecting Heterogeneous Treatment Effects
Results in Table 17 suggest that our dataset lacks sufficient statistical power to reliably detect race-specific heterogeneous treatment effects. This means that the absence of significant findings should not be interpreted as evidence of no heterogeneity. In our analysis, we focus on race-based subgroup analysis using the parameters from original MAAPS experiments: a total sample size of 745 students, a treated proportion of 52.6%, and a subgroup proportion of Black students of 47%. Under these real settings (Model 0), the minimum detectable effect (MDE) is 18.1 percentage points. This implies that our study is only powered to detect relatively large differences in graduation rates between different race-ethnicity subgroups, for example, an increase from 20% to 38.1% in the graduation outcome with 80% power at a 5% significance level.
Model # Sample Size Treated Proportion Sub-group Proportion Power Significance MDE
0 745 52.6% 47% 0.80 0.05 18.1%
1 1,000 52.6% 47% 0.80 0.05 15.6%
2.1 7,000 52.6% 20% 0.80 0.05 7.5%
2.2 7,000 52.6% 47% 0.80 0.05 6%
3 10,000 52.6% 47% 0.80 0.05 5%
Table 17: Of the additive minimal detectable effects (MDEs) from the power analysis on the synthetic dataset for which we specify population sample size, original treatment group size (52.6% for all experiments), sub-group proportion (based on a set covariate like race-ethnicity or assignment to particular MAAPS advisor), power level of 80%, and 5% significance.
Each model # (e.g., 1 vs. 2) indicates an index for a separate experiment with varying population sample sizes. Experiments with the same whole model number (e.g., 2.1 and 2.2) have the same total sample size but differ in the sub-group proportions.
Additionally, to explore the effects of increasing sample size while keeping subgroup and treatment proportions the same as in the original experiment, we simulate synthetic datasets with larger populations. As shown in Models 1 through 3 in Table 17, increasing the sample size from 1,000 to 10,000 students steadily reduces the MDE from 15.6% to 5%. Similarly, increasing the sub-group proportion reduces the MDE, as seen in Models 2.1 and 2.2, where raising the proportion from 20% to 47% lowers the MDE by approximately 1.5 percentage points.
A.6.7 Heterogeneous Treatment Effect Findings by Student Demographics and Assigned Advisor
We present the logistic regression estimates for the impact of advisor assignment on student graduation outcome in Table 18. The HTE estimates from this regression and the causal forest are summarized in Table 19 along with their BH-adjusted $p$-values. The results indicate that no single advisor is statistically more effective than another in terms of the measured outcome variables, as all advisor coefficients are small and nonsignificant. This suggests that while the advising styles vary, these differences do not necessarily translate into measurable impacts on student success outcomes. For Tables 18 and 19, we additionally visualize the estimates. Specifically, we plot the confidence intervals for the logistic regression interaction coefficients and causal forest HTE estimates in Figure 9 by advisor assignment on the student graduation outcome. To construct the interquartile ranges and confidence intervals for the logistic regression estimates, we used bootstrapping with 1,000 samples. For the causal forest HTEs, we plot the confidence intervals based on the effect intervals computed using the native effect_interval functionality. We observe a consistent pattern suggesting that Advisor 3 may be more effective in influencing student graduation outcomes, as indicated by both the causal forest analysis and the logistic regression findings.
We extend our analysis to examine race-specific HTEs for which logistic regression estimates are displayed in Table 20. As further shown in the summary Table 21, after correcting for multiple hypothesis testing, we find no statistically significant HTEs across all racial groups for the graduation outcome. Similar to the advisor analysis, for Tables 20 and 21, we plot the confidence intervals for both the logistic regression interaction coefficients and the causal forest HTE estimates in Figure 10. We construct the interquartile ranges and confidence intervals using the same approach as in advisor estimates.
Logit Regression (Graduation)
Variable Coef. Std. Err. $z$-value P$>|z|$ 95% Confidence Interval
Constant -0.5636 0.312 -1.804 0.285 [-1.176, 0.049]
Assigned Group 0.0978 0.390 0.251 0.962 [-0.667, 0.862]
Advisor 1 -0.2548 0.443 -0.576 0.847 [-1.122, 0.613]
Advisor 2 0.0204 0.424 0.048 0.962 [-0.810, 0.851]
Advisor 3 0.5109 0.416 1.227 0.495 [-0.305, 1.327]
Gender -0.6857 0.189 -3.625 0.001 [-1.056, -0.315]
ACT Composite 0.1805 0.093 1.947 0.155 [-0.001, 0.362]
EFC 0.0100 0.084 0.118 0.962 [-0.156, 0.176]
Hours Other 0.0866 0.088 0.982 0.587 [-0.086, 0.260]
Observations: 745, Pseudo R-squared: 0.02844, Log-Likelihood: -417.10, LLR $p$-value: 0.001946
Table 18: Logistic regression results estimating the effect of student assignment to a particular advisor on graduation outcomes. The model includes controls for gender, ACT scores, expected family contribution (EFC), and other academic hours applied toward the degree. Coefficient estimates, associated $z$-statistics, BH-adjusted $p$-values, and 95% confidence intervals are reported.
Advisor Logistic Regression Causal Forest
Coef1 $p$-value HTE $p$-value
Advisor 1 -0.2548 0.847 -0.0288 0.8547
Advisor 2 0.0204 0.962 -0.0163 0.7251
Advisor 3 0.5109 0.495 0.1049 0.4817
Table 19: Summary of heterogeneous treatment effects (HTEs) by advisor for graduation outcomes. We report logistic regression interaction coefficients $Coef1$, causal forest average $HTE$ estimates, and their BH-adjusted $p$-values.
Figure 9: Confidence intervals for the causal forest HTE estimates (left) and logistic regression interaction coefficients (right) by advisor on the graduation outcome. Each point marks the median estimate for a given race-ethnicity subgroup, with shaded boxes indicating the interquartile range (25th–75th percentiles) and vertical lines representing the 95% confidence intervals. Advisor 3 exhibits a higher positive median interaction effect and HTE estimate compared to other subgroups.
Logit Regression (Graduation)
Variable Coef. Std. Err. $z$-value P$>|z|$ 95% Confidence Interval
Constant -0.5636 0.312 -1.804 0.285 [-1.176, 0.049]
Assigned Group 0.0948 0.411 0.230 0.917 [-0.712, 0.901]
Is_Hispanic -0.0568 0.417 -0.136 0.917 [-0.873, 0.760]
T x Hispanic -0.3076 0.577 -0.533 0.891 [-1.438, 0.823]
Is_Asian -0.5436 0.435 -1.250 0.507 [-1.396, 0.309]
T x Asian 0.1557 0.565 0.276 0.917 [-0.952, 1.263]
Is_Black -0.5554 0.370 -1.500 0.401 [-1.281, 0.170]
T x Black 0.3425 0.485 0.707 0.822 [-0.607, 1.292]
Gender -0.6257 0.189 -3.313 0.011 [-0.996, -0.255]
ACT Composite 0.1848 0.096 1.922 0.285 [-0.004, 0.373]
EFC -0.0089 0.085 -0.104 0.917 [-0.176, 0.159]
Hours Other 0.0719 0.089 0.809 0.822 [-0.102, 0.246]
Observations: 745, Pseudo R-squared: 0.02656, Log-Likelihood: -417.91, LLR $p$-value: 0.01884
Table 20: Logistic regression results estimating treatment effect heterogeneity by race-ethnicity on graduation outcomes. This model includes interaction terms between treatment status and student race categories (Black, Hispanic, Asian), along with controls for gender, ACT score, EFC, and additional college credit hours earned. Coefficient columns report estimates with associated $z$-values, BH-adjusted $p$-values, and 95% confidence intervals.
Race Logistic Regression Causal Forest
Coef1 $p$-val Coef2 $p$-val HTE $p$-val
Black 0.3425 0.822 0.4373 0.2575 0.1087 0.5426
Hispanic -0.3076 0.891 -0.2128 0.5980 0.0513 0.7237
Asian 0.1557 0.917 0.2506 0.5981 0.0619 0.6714
Table 21: Summary of heterogeneous treatment effects (HTEs) by race-ethnicity for graduation outcomes. We report logistic regression interaction coefficients $Coef1$, combined treatment effect coefficients $Coef2$, causal forest average $HTE$ estimates, and their BH-adjusted $p$-values.
Coef1 represents the coefficient of the interaction term ($T_{i}\times\text{Race\_Ethnicity}_{i}$) in Equation (2). Coef2 is computed as $(T_{i}+T_{i}\times\text{Race\_Ethnicity}_{i})$. The associated $p$-value for Coef2 is the BH-adjusted significance level.
Figure 10: Confidence intervals for the causal forest HTE estimates (left) and logistic regression interaction coefficients (right) by race-ethnicity for the student graduation outcome. Each point marks the median estimate for a given race-ethnicity subgroup, with shaded boxes indicating the interquartile range (25th–75th percentiles) and vertical lines representing the 95% confidence intervals. Black students exhibit a higher positive median interaction effect and HTE estimate compared to other subgroups.
A.6.8 Advisor Action Prediction Analysis and Findings.
For each of the five significant interventions identified in §5.2.1, we estimate two logistic regression models for each academic semester, restricted to nontransfer students who were actively enrolled during that term. The outcome variable $A_{i}$ in each model is binary, indicating whether an advisor applied a particular intervention during a meeting with $i$-th student. Formally, for a given intervention, $A_{i}=1$ if the advisor applied it during the meeting and $A_{i}=0$ otherwise. Each student may contribute multiple observations per semester if they attended multiple advising meetings. For each of these meetings, advisor actions are predicted separately. The input features include the full set of baseline student characteristics (see Appendix A.1.1), and all categorical variables, including race-ethnicity and gender, are one-hot encoded. The first Model (10) incorporates the race-ethnicity variable, while the second Model (11) omits it:
$\text{logit}\left(\mathbb{P}(A_{i}=1)\right)=\beta_{0}+\beta_{1}T_{i}+\theta_{i}\text{Race}_{i}+\sum_{k=1}^{K}\gamma_{k}X_{k,i}+\epsilon_{i}$ (10)
$\text{logit}\left(\mathbb{P}(A_{i}=1)\right)=\beta_{0}+\beta_{1}T_{i}+\sum_{k=1}^{K}\gamma_{k}X_{k,i}+\epsilon_{i}$ (11)
Here, $T_{i}$ is the treatment indicator, $Race_{i}$ is the one-hot encoded race-ethnicity variable, and $X_{k,i}$ denotes the remaining student baseline features as listed in Appendix A.1.1. We fit separate regressions for each intervention per academic semester. AUC scores for advisor action predictions for each intervention are presented in Table 22, where we observe no numerical difference in the predictive performance between two models across all semesters.
Intervention 2016-Fa 2017-Sp 2017-Su 2017-Fa 2018-Sp 2018-Su 2018-Fa 2019-Sp 2019-Su 2019-Fa 2020-Sp
Intervention 2 0.445/0.445 0.590/0.590 0.473/0.473 0.511/0.511 0.564/0.564 0.246/0.246 0.545/0.545 0.510/0.510 0.638/0.638 0.553/0.553 0.553/0.553
Intervention 4 0.607/0.607 0.566/0.566 0.128/0.128 0.391/0.391 0.438/0.438 0.318/0.318 0.558/0.558 0.477/0.477 0.496/0.496 0.515/0.515 0.637/0.637
Intervention 16 0.534/0.534 0.591/0.591 0.566/0.566 0.518/0.518 0.499/0.499 0.563/0.563 0.589/0.589 0.609/0.609 -/- -/- -/-
Intervention 17 0.543/0.543 0.511/0.511 0.559/0.559 0.419/0.419 0.478/0.478 0.636/0.636 0.483/0.483 0.472/0.472 0.444/0.444 0.270/0.270 0.564/0.564
Intervention 20 0.512/0.512 0.573/0.573 0.504/0.504 0.555/0.555 0.557/0.557 0.509/0.509 0.563/0.563 0.424/0.424 0.513/0.513 0.487/0.487 -/-
Table 22: AUC scores for logistic regression classifiers predicting advisor intervention decisions for five significant interventions, across academic semesters. Each cell shows performance from two models: Model 1 (includes race-ethnicity) / Model 2 (excludes race-ethnicity).
For the missing entries, the action prediction is undefined due to the presence of only one class for that intervention.